### v1

# Inertial Magnetic SLAM Systems Using Low-Cost Sensors

Abstract
Spatially inhomogeneous magnetic fields offer a valuable, non-visual information source for positioning. Among systems leveraging this, magnetic field-based simultaneous localization and mapping (SLAM)systems are particularly attractive because they can provide positioning information and build a magnetic field map on the fly. Moreover, they have bounded error within mapped regions. However, state-of-the-art methods typically require low-drift odometry data provided by visual odometry or a wheel encoder, etc. This is because these systems need to minimize/reduce positioning errors while exploring, which happens when they are in unmapped regions. To address these limitations, this work proposes a loosely coupled and a tightly coupled inertial magnetic SLAM (IM-SLAM)system. The proposed systems use commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer. The use of non-visual data provides a significant advantage over visual-based systems, making it robust to low-visibility conditions. Both systems employ state-space representations, and magnetic field models on different scales. The difference lies in how they use a local and global magnetic field model. The loosely coupled system uses these models separately in two state-space models, while the tightly coupled system integrates them into one state-space model. Experiment results show that the tightly coupled IM-SLAMsystem achieves lower positioning errors than the loosely coupled system in most scenarios, with typical errors on the order of meters per 100 meters traveled. These results demonstrate the feasiblity of developing a full 3D IM-SLAMsystems using low-cost sensors and the potential of applying these systems in emergency response scenarios such as mine/fire rescue.
{IEEEkeywords}
inertial navigation, SLAM, indoor positioning, magnetic field, low-cost sensors.
[…]
1 Introduction
[…]
Positioning and navigation services are essential for various applications, including scientific research, emergency response, and military operations. These services have historically relied on the global navigation satellite system (GNSS). However, GNSS-denied environments, such as deep sea, indoor environments, and battlefields, where GNSS signals are unavailable or unreliable, pose significant challenges for accurate and reliable positioning. Therefore, developing robust and precise positioning systems that can operate effectively in GNSS-denied environments is of paramount importance.
A common approach to address the challenges in positioning in GNSS-denied environments is to use additional sensors, such as cameras, Light Detection and Ranging (LiDAR)sensors, and inertial measurement units (IMUs), to complement the GNSS. For example, visual and LiDARSimultaneous localization and mapping (SLAM) systems [1, 2] have been widely used for indoor positioning. However, these systems have their own limitations. Visual SLAM systems are sensitive to lighting conditions and may fail in low-light or featureless environments [3], and LiDAR SLAM systems can be expensive and may function poorly in environments with limited geometric features [4]. Thus, it is necessary to exploit other types of environmental features to achieve robust indoor SLAM.
Indoor magnetic fields have been recognized as a promising source of environmental features for SLAM [5]. Because of the ubiquitous presence of ferromagnetic materials in buildings, the indoor magnetic field exhibits spatial variations, which can be used as location fingerprints. As is shown in Fig. 1, the magnetic field strength inside a building can vary significantly over short distances, making it a useful feature for SLAMapplications. Moreover, magnetic field measurements can be obtained using low-cost magnetometers, which are commonly found in smartphones and other portable devices. This makes magnetic field SLAMsystems cost-effective and accessible.
Previous works have shown that magnetic field SLAMsystems can achieve meter-level accuracy in indoor environments [6, 7, 8, 9, 10]. However, most of these systems rely on low-drift odometry data, such as that from wheel odometry or visual odometry, to provide motion constraints, or they assume planar 2D motion, which limits their availability and applicability in real-world scenarios. Furthermore, these systems use a single magnetometer, which makes SLAMmore challenging because a measurement of the magnetic field at a single point provides limited information about the spatial variations in the magnetic field. To address these issues, in a recent work [11], a magnetic field-aided inertial navigation system was developed. This system uses a low-cost IMUand a magnetometer array consisting of 30 magnetometers, which allows snapshots of magnetic field over a planar sensor board to be taken. This system can provide low-drift inertial navigation that has the potential to be used, instead of wheel odometry or visual odometry, in a magnetic-field SLAM system. Inspired by this work, we seek to answer the following research questions:
Figure 1: Illustration of the magnetic-field magnitude variations, represented by different colors, near the floor in a room of $12\times 12$ square meters.
  • Can a 3D inertial magnetic SLAM (IM-SLAM) system be built and provide meter-level positioning using only a low-cost IMU, barometer, and magnetometer array?
  • What is a suitable sensor fusion strategy? How does the performance of a tightly-coupled IM-SLAMsystem compare to a loosely-coupled IM-SLAMsystems?
  • How does the barometer and the quality of the inertial sensor measurements affect the overall performance of the SLAM system?
To that end, we propose two IM-SLAMsystems: a loosely coupled and a tightly coupled system. Different from previous magnetic field SLAMsystems, the proposed systems do not require odometric information from visual or wheel-encoder sensors. This makes the systems much more adaptable to various environments where traditional odometry might fail.
1.1 Related Work
Many methods for magnetic field-based indoor positioning and navigation have been proposed. They are generally classified into two categories: infrastructure-based methods [12] and infrastructure-free methods [13, 14]. The former generally requires some magnetic source to be installed in the environment, such as magnetic coils or magnets. The latter, on the other hand, exploits the ambient magnetic field as a location fingerprint. The infrastructure-free methods are more appealing because they are easier to deploy and maintain. Therefore, our focus is on the infrastructure-free methods.
The infrastructure-free methods can be further classified into two categories: fingerprinting-based methods[15, 16, 17, 18, 14, 19] and model-based methods [20, 21, 11, 13, 7, 22, 10, 6]. The fingerprinting-based methods are based upon the assumption that the magnetic fields at different locations are distinct. These methods generally rely on matching real-time magnetic field measurements to a pre-established magnetic field map to estimate location, or to a stored measurement history to recognize previously visited locations for loop closure. The model-based methods, on the other hand, seek to use a mathematical model to describe magnetic fields measurements. These methods can use the magnetic measurements to update their magnetic field model (build a magnetic field map) and estimate pose change as the device moves through the environment. There are pros and cons for both methods. The fingerprinting-based methods can achieve high accuracy if magnetic field matching is successful, but they require carefully designed feature extraction and matching algorithms. Furthermore, erroneous matching can lead to significant localization errors. In contrast, the model-based methods are easier to analyze from a theoretical perspective, but they can be sensitive to the choice of the magnetic field model and the tuning of the model parameters.
The proposed IM-SLAMsystems are model-based. They were made possible by the development in magnetic field odometry and previous works on magnetic field modeling. In [20], a polynomial model was proposed to locally approximate the magnetic field, which was later extended in [21] to estimate pose changes of a magnetometer array between two time instants. The very same polynomial model was also used in [11] to build a magnetic field aided inertial navigation system (MAINS), which effectively reduced the positioning drift of the inertial navigation system by 2 orders of magnitude. This work constitutes one of the building blocks of the proposed SLAM systems. The other building block comes from the work in global magnetic field modeling. Originally proposed in [23], the reduced rank Gaussian process (GP)model was shown to be effective and computationally efficient in modeling the indoor magnetic field on a global scale. The GP model was later used in [24, 6, 7] to build magnetic field SLAM systems.
1.2 Contributions
The contributions of this work are as follows:
  • A loosely-coupled and tightly-coupled IM-SLAMsystems using low-cost sensors are proposed.
  • An extensive evaluation is conducted on real-world datasets, encompassing challenging indoor environments with multi-floor structures.
  • The implementation of the proposed systems and the datasets used for evaluation are made publicly available to provide a reproducible benchmark.
Reproducible research: Both the datasets and algorithms used in this study are available at https://github.com/Huang-Chuan/IM-SLAM.
2 Preliminaries
As magnetic field modeling is essential for magnetic field SLAM, we briefly review two commonly used magnetic field models in this section; the first model is for modeling the magnetic field at local small scales, and the second is at global large scales.
Let us start with magnetostatics. Assuming that there is no time-varying electric field and current density in the region of interest, the magnetic field fulfills the curl-free condition [25, p. 180]
$\nabla\times M(r)=0$ (1)
where $r\in\mathbb{R}^{3}$ denotes the position in the Cartesian coordinates. This permits the expression of $M(r)$ as the gradient of a scalar potential function $\phi(r):\mathbb{R}^{3}\rightarrow\mathbb{R}$, i.e.,
$M(r)=\nabla\phi(r).$ (2)
Both of the two magnetic field models presented next are based on (2).
2.1 Local Magnetic Field Model
A local magnetic field model refers to a model that captures the magnetic field in a small region of space where the field can be well approximated by a simple function. The polynomial model proposed in [20] is a model of this type. It provides a compact local representation of the magnetic field. In this model, the scalar potential function is assumed to be an $(n+1)$-degree polynomial function, i.e.,
$\phi(r;c)=\;\sum_{\begin{subarray}{c}i,j,k\geq 0\\ i+j+k\leq n+1\end{subarray}}c_{i,j,k}\,r_{x}^{i}r_{y}^{j}r_{z}^{k}.$ (3)
Here $c_{i,j,k}\in\mathbb{R}$ denotes the coefficient and $r\triangleq[r_{x}\;r_{y}\;r_{z}]^{\top}$. Furthermore, the coefficients of the resulting magnetic field model $M(r;c)$ from applying (2) are further constrained by the divergence-free condition [25, p. 180]
$\nabla\cdot M(r;c)=0.$ (4)
It leads to a $n$-th degree polynomial magnetic field model [20]
$M(r;\theta)=\Phi(r)\theta$ (5)
where $\Phi(r)\in\mathbb{R}^{3\times p}$ denotes the regressor matrix, $\theta\in\mathbb{R}^{p\times 1}$ denotes the coefficients, with $p=n^{2}+4n+3$.
It has been empirically verified in [11] that a first-degree polynomial magnetic field model can capture indoor magnetic fields on a planar magnetometer array of size $345\,\text{mm}\times 245\,\text{mm}$. The regressor matrix in this case is
$\Phi(r)=\begin{bmatrix}1&0&0&0&0&r_{z}&r_{y}&2r_{x}\\ 0&1&0&r_{z}&2r_{y}&0&r_{x}&0\\ 0&0&1&r_{y}&-2r_{z}&r_{x}&0&-2r_{z}\end{bmatrix}$ (6a)
and
$\theta=\left[\theta(1)\;\theta(2)\;\cdots\;\theta(8)\right]^{\top}.$ (6b)
2.2 Global Magnetic Field Model
A global magnetic field model refers to a model that describes the magnetic field across the entire environment, such as a building or a large lab. The approximated GPmodel proposed in [23] allows for efficient regression with much less computational complexity than the full GPmodel. The development of the magnetic field model starts with modeling the scalar magnetic potential function $\phi(r)$ as a GP
$\phi(r)\sim\mathcal{GP}\left(0,\kappa_{\text{lin.}}(r,r^{\prime})+\kappa_{\text{SE}}(r,r^{\prime})\right)$ (7a)
where
$\kappa_{\text{lin.}}(r,r^{\prime})=\sigma_{\text{lin.}}^{2}r^{\top}r^{\prime}$ (7b)
and
$\kappa_{\text{SE}}(r,r^{\prime})=\sigma_{\text{SE}}^{2}\text{exp}\left(-\frac{\lVert r-r^{\prime}\rVert^{2}}{2l^{2}_{\text{SE}}}\right).$ (7c)
Here $\sigma_{\text{lin.}}$ denotes the magnitude scale of the linear kernel $\kappa_{\text{lin.}}$, $\sigma_{\text{SE}}$ and $l_{\text{SE}}$ denote the magnitude scale and the characteristic length scale of the square exponential kernel $\kappa_{\text{SE}}$, respectively. To reduce the computation complexity, the square exponential kernel is approximated by
$\kappa_{\text{SE}}(r,r^{\prime})\approx\sum_{j=1}^{m}S_{\text{SE}}(\lambda_{j})\psi_{j}(r)\psi_{j}(r^{\prime})$ (8a)
where $m$ is the number of basic functions and $S_{\text{SE}}(\cdot)$ denotes the spectral density of $\kappa_{\text{SE}}$. Furthermore, $\psi_{j}(\cdot)$ and $\lambda_{j}$ denote the $j$-th eigenfunction and eigenvalue of the Laplace operator in the domain of a cuboid ($L_{x},L_{x}]\times[-L_{y},L_{y}]\times[-L_{z},L_{z}]$) with Dirichlet boundary conditions, respectively. They are given by
$\psi_{j}(r)$ $=\prod_{d=x,y,z}\frac{1}{\sqrt{L_{d}}}\sin\left(\frac{\pi n_{j,d}(r_{d}+L_{d})}{2L_{d}}\right)$ (8b)
$\lambda_{j}^{2}$ $=\sum_{d=x,y,z}\left(\frac{\pi n_{j,d}}{2L_{d}}\right)^{2}$ (8c)
respectively, where $n_{j,d}\in\mathbb{Z}^{+}$. Combining (7a) and (8a) and relating to the weight space view [26, Ch. 2], the scalar potential function is then approximated by
$\phi(r;\eta)\approx\Psi(r)\eta,\quad\eta\in\mathbb{R}^{3+m}$ (9a)
where
$\Psi(r)=[r^{\top}\;\psi_{1}(r)\;\psi_{2}(r)\;\cdots\;\psi_{m}(r)]$ (9b)
$\eta\sim\mathcal{N}(0,\Lambda)$ (9c)
and
$\Lambda=\text{diag}\left(\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},S_{\text{SE}}(\lambda_{1}),S_{\text{SE}}(\lambda_{2}),\cdots,S_{\text{SE}}(\lambda_{m})\right).$ (9d)
Here $\mathcal{N}(0,\Lambda)$ denotes a Gaussian distributed random variable with mean 0 and covariance matrix $\Lambda$. The global magnetic field model is obtained by applying (2) to the approximated scalar potential function, i.e.,
$M(r;\eta)=\nabla\phi(r;\eta)\approx\nabla\Psi(r)\eta.$ (10)
2.3 Discusssion
The polynomial model presented in Section 2.1 is well-suited for local modeling, due to its simplicity and efficiency. The first-degree polynomial model was used in [11] to model the magnetic field within the small area covered by the magnetometer array at each time step (see Fig. 2). In this model, both the spatial coordinates and the magnetic field are expressed in the body frame of the magnetometer array.
On the other hand, the approximated GPmodel presented in Section 2.2 is more appropriate for global modeling, as the sinusoidal basis functions can capture the magnetic field variations across a larger area. It is more computationally demanding than the first-degree polynomial model. Typically, several hundreds to thousands of basis functions are used to represent the magnetic field across an area of several hundred square meters [7]. Since this model aims to describe the field over a global region, both the spatial coordinates and the magnetic field are expressed in the navigation frame.
Figure 2: Coordinate frames used by the local and global magnetic field models. The local magnetic field model at each time step aligns with the body frame of the magnetometer array, e.g., $\mathcal{B}_{1},\mathcal{B}_{2}$ and $\mathcal{B}_{3}$, while the global magnetic field model uses the navigation frame denoted by $O$ in the middle of the plot.
3 Magnetic Field-Based Positioning Systems
In this section, we briefly review the system models in two magnetic field-based positioning systems: the first is the magnetic field SLAM system proposed in [7], and the second is the MAINSproposed in [11]. These systems, illustrated in Fig. 3a and Fig. 3b, are the foundations of the proposed IM-SLAMsystems.
Figure 3: (a) An overview of the magnetic field SLAM system [7]. The system performs dead reckoning using visual odometry data and uses a global magnetic field model to register a single magnetometer’s measurements, which are used to correct position drift. (b) An overview of the MAINS. The system performs inertial navigation using IMUmeasurements and uses a local magnetic field model to register a magnetometer array’s measurements, which are used to correct velocity estimates.
3.1 The Magnetic Field SLAM System
The state-of-the-art magnetic field SLAM system [7] uses low-drift visual odometry and magnetic field measurements from a single magnetometer. The basic idea is to perform dead-reckoning using the visual odometry data and use magnetic field measurements to correct the accumulated position drift.
The state vector of the magnetic field SLAM at time step $k$ is defined as
$x_{k}^{\text{\tiny{SLAM}}}\triangleq[p_{k}^{\top}\;q_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (11)
where $p_{k}\in\mathbb{R}^{3}$ denotes the position, $q_{k}\in S^{3}$ denotes the orientation in quaternion representation, and $\eta_{k}\in\mathbb{R}^{3+m}$ denotes the coefficients of the approximated GPmagnetic field model.
The system model of the magnetic field SLAM is
$x_{k+1}^{\text{\tiny{SLAM}}}$ $=f^{\text{\tiny{SLAM}}}(x_{k}^{\text{\tiny{SLAM}}},\tilde{u}_{k}^{o})$ (12a)
$y_{k}^{(0)}$ $=h^{\text{\tiny{SLAM}}}(x_{k}^{\text{\tiny{SLAM}}})+e_{k}^{(0)}$ (12b)
where
$f^{\text{\tiny{SLAM}}}(x_{k}^{\text{\tiny{SLAM}}},\tilde{u}_{k}^{o})=\begin{bmatrix}p_{k}+\Delta p_{k}\\ q_{k}\otimes\Delta q_{k}\\ \eta_{k}\end{bmatrix}$ (12c)
and
$h^{\text{\tiny{SLAM}}}(x_{k}^{\text{\tiny{SLAM}}})=R_{k}^{\top}\nabla\Psi\big(p_{k}\big)\eta_{k}.$ (12d)
Here $\tilde{u}_{k}^{o}=[\Delta p_{k}^{\top}\;\Delta q_{k}^{\top}]^{\top}$ denotes the visual odometry data, where $\Delta p_{k}\in\mathbb{R}^{3}$ and $\Delta q_{k}\in S^{3}$ denote the change in position and orientation, respectively. Furthermore, $y_{k}^{(0)}\in\mathbb{R}^{3}$ and $e_{k}^{(0)}\in\mathbb{R}^{3}$ are the magnetometer measurements and the Gaussian white measurement noise, respectively. In addition, $\otimes$ denotes quaternion multiplication, and $R_{k}\in SO(3)$ denotes the rotation matrix corresponding to the quaternion $q_{k}$.
3.2 The Magnetic Field-Aided Inertial Navigation System
The MAINSproposed in [11] uses a low-cost IMUand magnetometer array (see Fig. 4) to provide low-drift inertial navigation. The basic idea is to use inertial navigation system (INS)to estimate poses, and magnetic field measurements to correct velocity estimates.
The state vector of MAINS at time step $k$ is defined as
$x_{k}^{\text{\tiny{mains}}}\triangleq[{x_{k}^{\text{\tiny{ins}}}}^{\top}\;\theta_{k}^{\top}]^{\top}$ (13a)
$x_{k}^{\text{\tiny{ins}}}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}]^{\top}$ (13b)
where $x_{k}^{\text{\tiny{ins}}}$ denotes the inertial navigation states. Furthermore, $v_{k}\in\mathbb{R}^{3}$ denotes the velocity, $b_{a,k}\in\mathbb{R}^{3}$ and $b_{g,k}\in\mathbb{R}^{3}$ denote the accelerometer and gyroscope biases, respectively, and $\theta_{k}\in\mathbb{R}^{8}$ denotes the coefficients of the first degree polynomial magnetic field model in Section 2.1. The dynamics of the state vector and the measurement model are
$x_{k+1}^{\text{\tiny{mains}}}$ $=f^{\text{\tiny{mains}}}(x_{k}^{\text{\tiny{mains}}},\tilde{u}_{k},w_{k})$ (14a)
$y_{k}$ $=H^{\text{\tiny{mains}}}x_{k}^{\text{\tiny{mains}}}+e_{k}$ (14b)
where
$f^{\text{\tiny{mains}}}(x_{k}^{\text{\tiny{mains}}},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{ins}}}(x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\text{\tiny{ins}}}_{k})\\ f^{\theta}(\theta_{k},x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\theta}_{k})\end{bmatrix}$ (14c)
and
$H^{\text{\tiny{mains}}}=\begin{bmatrix}0_{3\times 16}&\Phi(r^{(1)})\\ \vdots&\vdots\\ 0_{3\times 16}&\Phi(r^{(N)})\\ \end{bmatrix}$. (14d)
Here $f^{\text{\tiny{ins}}}(\cdot)$ denotes the inertial subsystem dynamics, $f^{\theta}(\cdot)$ denotes the magnetic subsystem function. The exact form of these functions is presented in [11]. Moreover, $\tilde{u}_{k}=[\tilde{a}_{k}^{\top}\;\tilde{\omega}_{k}^{\top}]^{\top}$ denotes the IMU measurements, where $\tilde{a}_{k}\in\mathbb{R}^{3}$ and $\tilde{\omega}_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity, respectively. Furthermore, $w_{k}=[(w_{k}^{\text{\tiny{ins}}})^{\top}\;(w_{k}^{\theta})^{\top}]^{\top}$ denotes the process noise, where $w_{k}^{\text{\tiny{ins}}}\in\mathbb{R}^{12}$ and $w_{k}^{\theta}\in\mathbb{R}^{8}$ denote the process noise of the inertial and magnetic subsystems, respectively. The measurement vector $y_{k}\in\mathbb{R}^{3N}$ denotes the magnetometer measurements from an array of $N$ magnetometers, where $r^{(i)}$ denotes the position of the $i$-th magnetometer in the array expressed in the body frame. $\Phi(\cdot)$ denotes the regressor matrix in (6a). Finally, $e_{k}\in\mathbb{R}^{3N}$ denotes the white Gaussian measurement noise.
Figure 4: The sensor board used in the magnetic field-aided inertial navigation system. It has 30 PNI RM3100 magnetometers and an Osmium MIMU 4844 IMU mounted on the bottom side. (a) Front side of the sensor board showing the magnetometer array. (b) Back side of the sensor board showing the IMU.
3.3 Summary
Both the magnetic field SLAM system and the MAINShave their own advantages and limitations. The magnetic field SLAM system has a bounded positioning error in mapped regions. However, its reliance on the low-drift visual odometry limits its usability. The MAINS, on the other hand, can provide low-drift inertial navigation, but it only maintains a local magnetic field map and, thus, cannot provide positioning with bounded error. In the next section, we propose a loosely-coupled magnetic field SLAM system that combines the advantages of both systems while mitigating their limitations.
4 The Loosely-Coupled IM-SLAMSystem
A loosely coupled IM-SLAM system can be regarded as a cascaded architecture composed of a modified MAINSand a modified magnetic field SLAM system, as illustrated in Fig. 5. The dependence on low-drift visual odometry data is eliminated by replacing it with odometry provided by the modified MAINSsystem. Furthermore, the measurement model in the magnetic field SLAM system is adapted to support magnetometer arrays, rather than a single magnetometer, as well as an additional barometer. The barometer supplies altitude information, enabling the SLAM system to achieve improved 3D trajectory estimation.
Figure 5: An overview of the loosely-coupled IM-SLAMsystem. The system is a cascade of a modified MAINS, an odometry data conversion module, and a modified magnetic field SLAM system.
4.1 Computing Odometry Data with the MAINS
The original MAINS [11] estimates the pose at the current timestep; therefore, it cannot be used directly to compute odometry data. One way to make the MAINS support odometry data computation is to augment the state vector of MAINSto include the complete state vector (or the pose) at the past and compute the difference between the pose at the pase and the current. A detailed explanation of this modification is presented next. The augmented state vector is defined as
$x_{k}^{a}$ $\triangleq[x_{k}^{\text{\tiny{mains}}\top}\;x_{k}^{\text{\tiny{past}}\top}]^{\top}$ (15a)
$x_{1}^{\text{\tiny{past}}}$ $\triangleq x_{1}^{\text{\tiny{mains}}}$ (15b)
$x_{k}^{\text{\tiny{past}}}$ $\triangleq f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}}),\quad k\geq 2$ (15c)
$f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})$ $=\begin{cases}x_{k-1}^{\text{\tiny{past}}}&nK+3\leq k\leq(n+1)K+1\\ x_{k-1}^{\text{\tiny{mains}}}&k=nK+2\end{cases}$ (15d)
$n$ $=0,1,2,\cdots$ (15e)
Here $K\in\mathbb{N}$ denotes the number of time steps within the time interval for which the odometry data is computed. The state-space model of the modified MAINSis then
$x_{k+1}^{a}=f^{a}(x_{k}^{a},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{mains}}}(x_{k}^{\text{\tiny{mains}}},\tilde{u}_{k},w_{k})\\ f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})\end{bmatrix}$ (16a)
$y_{k}$ $=H^{a}x_{k}^{a}+e_{k}$ (16b)
$H^{a}$ $=\begin{bmatrix}0_{3\times 16}&\Phi(r^{(1)})&0_{3\times 16}\\ \vdots&\vdots&\vdots\\ 0_{3\times 16}&\Phi(r^{(N)})&0_{3\times 16}\end{bmatrix}$. (16c)
The modified MAINS can compute odometric data between the time step $i=nK+1$ and $j=(n+1)K+1$ using the posterior state estimate at time step $j$ as
$\tilde{u}^{o}_{i,j}$ $=\begin{bmatrix}\Delta p_{i,j}\\ \Delta q_{i,j}\end{bmatrix}=\begin{bmatrix}\hat{p}_{j}-\hat{p}_{i}\\ \hat{q}_{i}^{\ast}\otimes\hat{q}_{j}\end{bmatrix}$ (17a)
where $(\hat{\cdot}_{i})$ and $(\hat{\cdot}_{j})$ denote the smoothed and filtered quantity at time $i$ and $j$, respectively, and $(\cdot)^{\ast}$ denotes conjugated quaternion. The covariance of the odometry data can be approximated by
$\text{Cov}(\tilde{u}_{i,j}^{o})\approx A_{i,j}\;\text{Cov}(\hat{p}_{i},\hat{p}_{j},\hat{q}_{i},\hat{q}_{j})\;A_{i,j}^{\top}$ (17b)
$A_{i,j}=\begin{bmatrix}I_{3}&-I_{3}&0_{3}&0_{3}\\ 0_{3}&0_{3}&I_{3}&-\hat{R}_{i}^{\top}\hat{R}_{j}\end{bmatrix}$ (17c)
where $\text{Cov}(\cdot)$ denotes the covariance. Note the covariance of the quaternion is defined in the tangent space of $S^{3}$, see [27, p. 54] for details.
4.2 The Modified Magnetic Field SLAM System
The modified magnetic field SLAM system uses the odometric data provided by the modified MAINS, magnetic field measurements from the sensor array, and barometer measurements. The state vector of this system and the dynamics remain the same as in Section 3.1. The measurement equation is modified to
$y_{k}=h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})+e_{k}$ (18a)
where
$h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})=\begin{bmatrix}R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(1)}\big)\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(2)}\big)\\ \vdots\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(N)}\big)\end{bmatrix}\eta_{k}.$ (18b)
Here $h^{\text{\tiny{array-slam}}}(\cdot)$ denotes the modified measurement function, which takes into account the multiple sensors and the geometric configuration of the sensor array. This measurement equation is augmented when barometric measurements are available, i.e.,
$y_{k}^{b}=p_{z,k}+e_{k}^{b}$ (19)
where $p_{z,k}\in\mathbb{R}$ denotes the altitude (the third component of the position vector), and $e_{k}^{b}\in\mathbb{R}$ denotes the barometer measurement noise, modeled as Gaussian white noise.
4.3 Summary
The loosely-coupled IM-SLAM system combines the advantages of both the magnetic field SLAM system and the MAINS. It is a self-contained system that can provide bounded-error positioning in mapped areas without the need for external low-drift odometry. However, there are two main limitations of this system. First, the magnetometer measurements are used twice: first in the modified MAINSand then in the modified magnetic field SLAM system. Second, the odometric data provided by the modified MAINSis correlated over time, which violates the assumption of uncorrelated odometric data in the magnetic field SLAM system. These two limitations can potentially degrade the performance of the loosely-coupled IM-SLAM system. To address these limitations, we propose a tightly-coupled IM-SLAM system in the next section.
5 The Tightly-Coupled IM-SLAMSystem
The tightly-coupled IM-SLAM system integrates the inertial, barometric, and magnetic sensors into a unified SLAM framework (see Fig. 6). In this framework, the state vector encompasses the full set of states, including the inertial states, local and global magnetic field model coefficients. The system dynamics and measurement equation are jointly formulated to leverage all available information to achieve better performance. In the following, the state-space model of this system is presented.
Figure 6: An overview of the tightly-coupled IM-SLAMsystem. This system consists of an inertial navigation system, a barometer model, and a local and global magnetic field model. These components are integrated into a unified SLAMframework.
5.1 State Dynamics
The state vector of the tightly-coupled SLAM system at time step $k$ is defined as
$x_{k}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}\;\theta_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (20)
The state transition model is
$x_{k+1}=f(x_{k},\tilde{u}_{k},w_{k})$ (21)
where
$f(x_{k},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{ins}}}(x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\text{\tiny{ins}}}_{k})\\ f^{\theta}(\theta_{k},x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\theta}_{k})\\ \eta_{k}\end{bmatrix}$. (22)
5.2 Measurement Equations
Since the state vector includes both the local and global magnetic field model coefficients, the magnetometer array’s measurements can be expressed in terms of both models. To achieve IM-SLAM, one can use the measurement equations in an alternating manner from either the MAINSor the magnetic field SLAM system. A natural choice is to use the computationally efficient measurement equation from the MAINSmost of the time, while occasionally using the more complex measurement equation from the magnetic field SLAM system. In this way, the system can benefit from the low-drift inertial navigation provided by the MAINSwhile still leveraging the global consistency of the magnetic field SLAM system. In the proposed tightly-coupled IM-SLAM, the measurement equation of the global model, which is based on (18), is further simplified to reduce computational complexity. The basic idea is that the magnetic field at any sensor’s location can be viewed as a superimposition of the magnetic field in the center of the array and some local magnetic field variation. This results in
$y_{k}=h^{m}(x_{k})+e_{k}$ (23a)
where
$h^{m}(x_{k})=\begin{bmatrix}R_{k}^{\top}\nabla\Psi\big(p_{k}\big)\\ R_{k}^{\top}\nabla\Psi\big(p_{k}\big)\\ \vdots\\ R_{k}^{\top}\nabla\Psi\big(p_{k}\big)\\ \end{bmatrix}\eta_{k}+\begin{bmatrix}\Phi(r^{(1)})-\Phi(0)\\ \Phi(r^{(2)})-\Phi(0)\\ \vdots\\ \Phi(r^{(N)})-\Phi(0)\end{bmatrix}\theta_{k}.$ (23b)
Here, the first term computes the magnetic field at the center of the magnetometer array by the approximated GPmodel, while the second term captures the local variations across the array modeled by the polynomial model. Note that for the first-degree polynomial model, the second term can also be rewritten as
$(\Phi(r^{(i)})-\Phi(0))\theta_{k}=\text{grad}(\theta_{k})r^{(i)}\\$ (24a)
where
$\text{grad}(\theta_{k})\triangleq\begin{bmatrix}2\theta_{k}(8)&\theta_{k}(7)&\theta_{k}(6)\\ \theta_{k}(7)&2\theta_{k}(5)&\theta_{k}(4)\\ \theta_{k}(6)&\theta_{k}(4)&-2\left(\theta_{k}(8)+\theta_{k}(5)\right)\\ \end{bmatrix}$. (24b)
Here, the right-hand side of (24a) is more compact and computationally efficient than the left-hand side. Furthermore, $\text{grad}(\theta_{k})\in\mathbb{R}^{3\times 3}$ has a clear physical interpretation, i.e., the gradient of the local magnetic field.
To sum up, there are two measurement equations for magnetometer array measurements: the equation used by the MAINSand the equation (23). They are used in an alternating manner: the former is employed at every update step except once $D$ samples, when the second model is used instead. That is,
$y_{k}=\begin{cases}h^{m}(x_{k})+e_{k}&\text{for}\;k=D,2D,\cdots\\ [H^{\text{\tiny{mains}}}\quad 0]x_{k}+e_{k}&\text{Otherwise}\end{cases}$ (25)
Lastly, similar to what was done for the loosely coupled IM-SLAMsystem, the measurement equation is augmented when barometer measurements are available.
6 Experimental Results
To evaluate the performance of the proposed loosely-coupled and tightly-coupled IM-SLAMsystems, we conducted experiments in an indoor environment using a custom-built sensor array, see Fig. 4. The sensor array is equipped with an IMU, a barometer, and an array of magnetometers.
In each experiment, a person carried the sensor array and started moving in a room equipped with a motion capture system for several tens of seconds. The person then walked out of the room and explored the surrounding area for several minutes before returning to the room. In total, three types of trajectories were recorded: a trajectory with long straight segments, a trajectory with square loops, and a trajectory with spiral upward segments (climbing a spiral staircase). The characteristics of the datasets are summarized in TABLE 1.
The IMUand magnetometers were sampled at 100 Hz, while the barometer was sampled at 10 Hz. When available, the motion capture system provided ground truth poses at 100 Hz. In the loosely-coupled IM-SLAM system, the odometry data was internally computed at $5$ Hz (K=20K\!=\!20). In the tightly-coupled SLAM system, the magnetic field model switching time was set to 1 second (D=100D\!=\!100), the length scale of the approximated GPmodel $l_{\text{SE}}$ was set to 0.7 meters. More details about the parameter settings can be found in the code repository.
The standard error-state extended Kalman filter (EKF)is used in all systems for state estimation; interested reader are referred to [27] for details.
Table 1: Information about the datasets
Data sequence LC-1 LC-2 LC-3 C-1 C-2 C-3 SS-1 SS-2 SS-3
Trajectory duration* (s) 239 237 234.5 208 107 118 192 181 194
[…]
Average height (m) 0.84 0.93 0.80 1.14 1.04 1.02 2.29 2.39 2.17
Maximum height difference (m) 1.68 1.68 1.59 1.76 1.76 1.85 4.53 4.36 4.28
[…]
* including the initial part of the trajectory in the motion capture area, $\dagger$ excluding the initial part of the trajectory in the motion capture area.
+ the difference between the maximum and minimum magnetic field strength measured by the magnetometer array during the trajectory.
LC: long corridor   C: corridor   SS: spiral staircase.
6.1 Comparison with the MAINS
In the first experiment, the loosely-coupled and tightly-coupled IM-SLAM systems were compared with MAINS. Note that MAINS was modified in the same way as in Section 3.1 to integrate barometer measurements.
All three systems used the pose from the motion capture system as additional measurements when the sensor array was in the room at the beginning of the trajectory. This is to stablize the IMU bias estimates, which is important for good positioning performance. During the later parts of the trajectory, the systems relied solely on measurements from the onboard sensors for positioning.
Three examples of the trajectories estimated by the three systems are shown in Fig. 7, Fig. 8, and Fig. 9, respectively. The quantitative results are summarized in TABLE 2 and TABLE 3, where the horizontal error, vertical error, and yaw error at the end of the trajectories are reported. The horizontal error is defined as the Euclidean distance between the estimated position and the ground truth position projected onto the horizontal plane, while the vertical error is defined as the absolute difference between the estimated altitude and the ground truth altitude. The yaw error is defined as the absolute difference between the estimated yaw angle and the ground truth yaw angle.
It can be seen from the figures that all systems can track the general shape of the trajectories. The noticeable difference between the tightly coupled IM-SLAMand other systems is that the estimated trajectories by the former have more jerky segments, which occur when the systems switch between the two measurement models; see (25). This phenomenon could be attributed to two different modeling of the magnetic field models: the model in the first row of (25) is parameterized with the position of the body frame in the navigation frame while the model in the second row is parameterized with the (local) positions of the magnetometers in the body frame. This implies that the former allows direct observation of the position while the latter does not [28]. The sudden jump observed on the estimated trajectory is a sign of gaining positional information when switching from the latter to the former model. Another observation is that the trajectories estimated by the SLAM systems during the long straight segments are more concentrated than those estimated by MAINS (see Fig. 7), which indicates that the SLAM systems can better mitigate the drift during these segments. Moreover, from TABLE 2 and TABLE 3, it can be seen that the tightly-coupled IM-SLAM system outperforms both MAINS and the loosely-coupled IM-SLAM system in most cases.
Table 2: Horizontal (Vertical) Error at the end of the Trajectories with a Barometer. Unit: meter
[…]
Long corridor-1 3.53 (0.29) 9.20 (0.91) 0.90 (0.22)
Long corridor-2 2.28 (0.17) 5.23 (1.01) 1.51 (0.16)
Long corridor-3 1.25 (0.13) 2.44 (0.42) 1.08 (0.10)
Corridor-1 5.22 (0.14) 10.46 (0.27) 2.57 (0.14)
Corridor-2 1.93 (0.09) 1.03 (0.77) 1.93 (0.13)
Corridor-3 1.98 (0.03) 1.76 (0.10) 1.78 (0.05)
Spiral staircase-1 3.55 (0.02) 5.22 (0.26) 1.79 (0.00)
Spiral staircase-2 3.59 (0.17) 3.45 (0.12) 2.43 (0.16)
Spiral staircase-3 8.37 (0.20) 9.13 (0.84) 3.20 (0.21)
*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
Table 3: Yaw Error at the end of the Trajectories with a Barometer. Unit: degree
[…]
Long corridor-1 1.45 1.59 0.28
Long corridor-2 0.79 1.68 2.34
Long corridor-3 0.29 1.30 1.65
Corridor-1 2.02 2.85 1.77
Corridor-2 3.00 1.38 1.05
Corridor-3 2.19 1.57 0.52
Spiral staircase-1 2.32 1.06 1.83
Spiral staircase-2 0.64 2.35 1.40
Spiral staircase-3 1.58 2.49 0.52
*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
Figure 7: Trajectories (Long corridor-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 8: Trajectories (Corridor-3) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
Figure 9: Trajectories (Spiral staircase-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system. The room where the motion capture system is located is marked with a blue rectangle frame.
6.2 Ablation Study on the Effect of Barometer and Inertial Sensor Quality
In the second part of the experiment, the effect of the barometric measurements and inertial sensor quality on the performance of the algorithms was investigated. For the first part, the barometer measurements were removed from all three systems, and the results are summarized in TABLE 4 and TABLE 5. Comparing TABLE 4 with TABLE 2, it can be seen that both the vertical error and horizontal error of all three systems increase slightly without the barometer measurements, but the overall performance remains unchanged. Furthermore, the yaw error of all three systems do not change significantly as seen from TABLE 5. This indicates that the IM-SLAM systems are not heavily reliant on the barometer measurements. However, the barometer measurements are expected to have a more significant impact if explicit loop closure detection is to be performed in the SLAM systems, since the altitude information can help to distinguish different floors in a multi-floor building. For the second part, the IMU measurements quality is varied by averaging the measurements from different numbers of IMU sensors in Osmium MIMU 4844 IMU. The box plots of the horizontal errors at the end of the trajectories vs. the number of IMU sensors are shown in Fig. 10, 11, and 12. It can be seen that the horizontal errors of the tightly-coupled IM-SLAM system are less sensitive to the number of IMU sensors compared to the other two systems, and it outperforms the other two algorithms on the datasets Long corridor and Corridor. On Spiral staircase datasets, the performance of the tightly-coupled IM-SLAM system improves as the number of IMU sensors increases, but it is not as good as MAINSwhen the number of IMU sensors is less than 4. We do not have a clear explanation for this phenomenon yet, but it could be related to the fact that the datasets Spiral staircase have more dynamic motions (e.g., turning and climbing) than the other two types of datasets. Further investigation is needed to fully understand this phenomenon. Lastly, the sudden increase in horizontal error when using 32 IMUsensors is due to exceptionally worse performance on one of the datasets (Spiral staircase-3), which is currently under investigation.
Table 4: Horizontal (Vertical) Error at the End of the Trajectories without a Barometer. Unit: meter
[…]
Long corridor-1 3.80 (1.01) 6.84 (6.75) 0.97 (0.83)
Long corridor-2 2.33 (1.16) 8.41 (1.21) 1.49 (0.11)
Long corridor-3 1.19 (1.21) 4.30 (0.89) 1.23 (0.43)
Corridor-1 5.35 (0.54) 11.05 (0.36) 2.45 (0.28)
Corridor-2 1.82 (0.44) 0.99 (0.95) 1.76 (0.32)
Corridor-3 2.02 (0.52) 1.80 (0.36) 1.89 (0.51)
Spiral staircase-1 3.00 (0.41) 6.90 (1.08) 1.74 (0.16)
Spiral staircase-2 3.39 (0.10) 4.33 (0.90) 2.51 (0.22)
Spiral staircase-3 8.22 (0.65) 8.48 (2.14) 4.13 (0.91)
*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
Table 5: Yaw Error at the End of the Trajectories without a Barometer. Unit: degree
[…]
Long corridor-1 0.87 2.85 0.29
Long corridor-2 1.24 2.17 2.31
Long corridor-3 0.23 1.76 2.00
Corridor-1 2.11 3.12 1.32
Corridor-2 2.90 1.37 0.90
Corridor-3 2.11 1.60 0.28
Spiral staircase-1 2.25 0.23 1.62
Spiral staircase-2 0.56 2.20 0.80
Spiral staircase-3 1.71 0.41 1.40
*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
Figure 10: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Long corridor datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
Figure 11: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Corridor datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
Figure 12: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Spiral staircase datasets. The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
7 Conclusion
In this paper, a loosely coupled and a tightly coupled IM-SLAMsystem were proposed. Both systems rely solely on commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer. The proposed systems have position errors on the order of meters per 100 meters traveled, which showcases the feasibility of developing magnetic field-based SLAMsystems with low-cost sensors. Experimental results demonstrate that the tightly coupled IM-SLAMsystem achieves better performance than the loosely coupled counterpart. This improvement arises because the tightly coupled system integrates the local and global magnetic field model within a unified framework, rather than employing them in a two-stage manner as in the loosely coupled design. Furthermore, the experiments show that incorporating a barometer primarily benefits IM-SLAMperformance in the vertical direction, and that the IMU quality can be degraded by approximately a factor of 32 without substantially affecting the overall accuracy. These results demonstrate the potential of applying the proposed systems in emergency response scenarios such as mine/fire rescue.

### v2

@@
 Spatially inhomogeneous magnetic fields offer a valuable, non-visual information source for positioning.
-Among systems leveraging this, magnetic field-based simultaneous localization and mapping (SLAM)systems are particularly attractive because they can provide positioning information and build a magnetic field map on the fly.
-Moreover, they have bounded error within mapped regions.
-However, state-of-the-art methods typically require low-drift odometry data provided by visual odometry or a wheel encoder, etc.
-This is because these systems need to minimize/reduce positioning errors while exploring, which happens when they are in unmapped regions.
-To address these limitations, this work proposes a loosely coupled and a tightly coupled inertial magnetic SLAM (IM-SLAM)system.
-The proposed systems use commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer.
-The use of non-visual data provides a significant advantage over visual-based systems, making it robust to low-visibility conditions.
-Both systems employ state-space representations, and magnetic field models on different scales.
-The difference lies in how they use a local and global magnetic field model.
-The loosely coupled system uses these models separately in two state-space models, while the tightly coupled system integrates them into one state-space model.
-Experiment results show that the tightly coupled IM-SLAMsystem achieves lower positioning errors than the loosely coupled system in most scenarios, with typical errors on the order of meters per 100 meters traveled.
-These results demonstrate the feasiblity of developing a full 3D IM-SLAMsystems using low-cost sensors and the potential of applying these systems in emergency response scenarios such as mine/fire rescue.
-{IEEEkeywords}
+Among systems leveraging this, magnetic field-based simultaneous localization and mapping (SLAM) systems are particularly attractive.
+These systems execute positioning and magnetic field mapping tasks simultaneously, and they have bounded positioning error within previously visited regions.
+However, state-of-the-art magnetic-field SLAM methods typically require low-drift odometry data provided by visual odometry, a wheel encoder, or pedestrian dead-reckoning technology.
+To address this limitation, this work proposes loosely coupled and tightly coupled inertial magnetic SLAM (IM-SLAM) systems, which use only low-cost sensors: an inertial measurement unit (IMU), 30 magnetometers, and a barometer.
+Both systems are based on a magnetic-field-aided inertial navigation system (INS) and use error-state Kalman filters for state estimation.
+The key difference between the two systems is whether the navigation state estimation is done in one or two steps.
+These systems are evaluated in real-world indoor environments with multi-floor structures.
+The results of the experiment show that the tightly coupled inertial magnetic SLAM (IM-SLAM) system achieves lower positioning errors than the loosely coupled system in most scenarios, with typical errors on the order of meters per 100 meters traveled.
+These results demonstrate the feasibility of developing a full 3D IM-SLAM system using low-cost sensors.
+A potential application of the proposed systems is for the positioning of first responders.
+keywords:
 inertial navigation, SLAM, indoor positioning, magnetic field, low-cost sensors.
-[…]
 1 Introduction
@@
 Positioning and navigation services are essential for various applications, including scientific research, emergency response, and military operations.
-These services have historically relied on the global navigation satellite system (GNSS).
-However, GNSS-denied environments, such as deep sea, indoor environments, and battlefields, where GNSS signals are unavailable or unreliable, pose significant challenges for accurate and reliable positioning.
-Therefore, developing robust and precise positioning systems that can operate effectively in GNSS-denied environments is of paramount importance.
-A common approach to address the challenges in positioning in GNSS-denied environments is to use additional sensors, such as cameras, Light Detection and Ranging (LiDAR)sensors, and inertial measurement units (IMUs), to complement the GNSS.
-For example, visual and LiDARSimultaneous localization and mapping (SLAM) systems [1, 2] have been widely used for indoor positioning.
+These services have historically relied on positioning information from global navigation satellite systems (GNSS).
+However, GNSS-denied environments, such as indoors or subsurface, where GNSS signals are unavailable or unreliable, pose significant challenges for accurate and reliable positioning.
+Therefore, developing robust and accurate positioning systems that can operate effectively in GNSS-denied environments is of paramount importance.
+A common approach to address the challenges of positioning in GNSS-denied environments is to use additional sensors, such as cameras, Light Detection and Ranging (LiDAR), and inertial measurement units (IMU s), to complement the GNSS.
+For example, visual and LiDAR SLAM systems [1, 2] have been widely used for indoor positioning.
 However, these systems have their own limitations.
@@
 Thus, it is necessary to exploit other types of environmental features to achieve robust indoor SLAM.
-Indoor magnetic fields have been recognized as a promising source of environmental features for SLAM [5].
-Because of the ubiquitous presence of ferromagnetic materials in buildings, the indoor magnetic field exhibits spatial variations, which can be used as location fingerprints.
-As is shown in Fig.
-1, the magnetic field strength inside a building can vary significantly over short distances, making it a useful feature for SLAMapplications.
-Moreover, magnetic field measurements can be obtained using low-cost magnetometers, which are commonly found in smartphones and other portable devices.
+Spatial variations of indoor magnetic fields have been recognized as a promising source of environmental features for SLAM [5].
+Because of the ubiquitous presence of ferromagnetic materials in buildings, the indoor magnetic field exhibits spatial variations over short distances.
+Moreover, indoor magnetic fields can be measured using low-cost magnetometers, which are commonly found in smartphones and other portable devices.
 This makes magnetic field SLAM systems cost-effective and accessible.
 Previous works have shown that magnetic field SLAM systems can achieve meter-level accuracy in indoor environments [6, 7, 8, 10, 9].
-However, most of these systems rely on low-drift odometry data, such as that from wheel odometry or visual odometry, to provide motion constraints, or they assume planar 2D motion, which limits their availability and applicability in real-world scenarios.
+However, most of these systems rely on low-drift odometry data, such as that from wheel odometry or visual odometry, to provide motion constraints, or they assume planar 2D motion, which limits their use in any type of practically useful positioning system.
 Furthermore, these systems use a single magnetometer, which makes SLAM more challenging because a measurement of the magnetic field at a single point provides limited information about the spatial variations in the magnetic field.
-To address these issues, in a recent work [11], a magnetic field-aided inertial navigation system was developed.
-This system uses a low-cost IMUand a magnetometer array consisting of 30 magnetometers, which allows snapshots of magnetic field over a planar sensor board to be taken.
+To address these issues, in a recent work [11], a magnetic-field-aided inertial navigation system was developed.
+This system uses a low-cost IMU and an array of 30 magnetometers, which allows snapshots of the magnetic field to be taken.
 This system can provide low-drift inertial navigation that has the potential to be used, instead of wheel odometry or visual odometry, in a magnetic-field SLAM system.
 Inspired by this work, we seek to answer the following research questions:
-Figure 1: Illustration of the magnetic-field magnitude variations, represented by different colors, near the floor in a room of $12\times 12$ square meters.
-  • Can a 3D inertial magnetic SLAM (IM-SLAM) system be built and provide meter-level positioning using only a low-cost IMU, barometer, and magnetometer array?
-  • What is a suitable sensor fusion strategy? How does the performance of a tightly-coupled IM-SLAMsystem compare to a loosely-coupled IM-SLAMsystems?
-  • How does the barometer and the quality of the inertial sensor measurements affect the overall performance of the SLAM system?
+  • Can a 3D inertial magnetic SLAM (IM-SLAM) system be built using only a low-cost IMU, barometer, and magnetometer array?
+  • How does the performance of a tightly coupled IM-SLAM system compare to a loosely coupled IM-SLAM system?
+  • How do the barometer and the quality of the inertial sensor measurements affect the overall performance of the SLAM system?
 To that end, we propose two IM-SLAM systems: a loosely coupled and a tightly coupled system.
 Different from previous magnetic field SLAM systems, the proposed systems do not require odometric information from visual or wheel-encoder sensors.
-This makes the systems much more adaptable to various environments where traditional odometry might fail.
+We evaluate the proposed systems on real-world datasets to answer the research questions.
 1.1 Related Work
-Many methods for magnetic field-based indoor positioning and navigation have been proposed.
+Many methods for magnetic-field-based indoor positioning have been proposed.
 They are generally classified into two categories: infrastructure-based methods [12] and infrastructure-free methods [13, 14].
@@
 The infrastructure-free methods are more appealing because they are easier to deploy and maintain.
-Therefore, our focus is on the infrastructure-free methods.
-The infrastructure-free methods can be further classified into two categories: fingerprinting-based methods[15, 16, 17, 18, 14, 19] and model-based methods [20, 21, 11, 13, 7, 22, 10, 6].
-The fingerprinting-based methods are based upon the assumption that the magnetic fields at different locations are distinct.
-These methods generally rely on matching real-time magnetic field measurements to a pre-established magnetic field map to estimate location, or to a stored measurement history to recognize previously visited locations for loop closure.
+The infrastructure-free methods can be further classified into two categories: fingerprint-based methods[15, 16, 17, 18, 14, 19] and model-based methods [20, 21, 11, 13, 7, 22, 9, 6].
+The fingerprint-based methods are based upon the assumption that the magnetic fields at different locations are distinct.
+These methods generally rely on matching magnetic field measurements to a pre-established magnetic field map to estimate position, or to a stored measurement history to recognize previously visited locations for loop closure.
 The model-based methods, on the other hand, seek to use a mathematical model to describe magnetic fields measurements.
@@
 There are pros and cons for both methods.
-The fingerprinting-based methods can achieve high accuracy if magnetic field matching is successful, but they require carefully designed feature extraction and matching algorithms.
-Furthermore, erroneous matching can lead to significant localization errors.
-In contrast, the model-based methods are easier to analyze from a theoretical perspective, but they can be sensitive to the choice of the magnetic field model and the tuning of the model parameters.
+The fingerprint-based methods can achieve high accuracy if magnetic field matching is successful, but they require carefully designed feature extraction and matching algorithms.
+Furthermore, erroneous matching can lead to significant positioning errors.
+In contrast, the model-based methods are easier to establish a theoretically sound framework for inference process, but they can be sensitive to the choice of the magnetic field model and the tuning of the model parameters.
 The proposed IM-SLAM systems are model-based.
-They were made possible by the development in magnetic field odometry and previous works on magnetic field modeling.
-In [20], a polynomial model was proposed to locally approximate the magnetic field, which was later extended in [21] to estimate pose changes of a magnetometer array between two time instants.
-The very same polynomial model was also used in [11] to build a magnetic field aided inertial navigation system (MAINS), which effectively reduced the positioning drift of the inertial navigation system by 2 orders of magnitude.
-This work constitutes one of the building blocks of the proposed SLAM systems.
+They were made possible by developments in magnetic field odometry and previous work on magnetic field modeling.
+In [20], a polynomial model was proposed to locally approximate the magnetic field, which was later extended in [21] to estimate pose changes of a magnetometer array.
+This polynomial model was also used in [11] to build a magnetic field aided inertial navigation system (MAINS), which effectively reduced the positioning drift of the inertial navigation system by 2 orders of magnitude.
+This work [11] constitutes one of the building blocks of the proposed SLAM systems.
 The other building block comes from the work in global magnetic field modeling.
@@
 The contributions of this work are as follows:
-  • A loosely-coupled and tightly-coupled IM-SLAMsystems using low-cost sensors are proposed.
-  • An extensive evaluation is conducted on real-world datasets, encompassing challenging indoor environments with multi-floor structures.
-  • The implementation of the proposed systems and the datasets used for evaluation are made publicly available to provide a reproducible benchmark.
+  • We propose a loosely coupled and tightly coupled 3D IM-SLAM system, which use only a low-cost IMU, a magnetometer array, and a barometer.
+These two system architectures serve as practical reference designs for practitioners.
+  • We conduct an evaluation of the proposed systems in terms of their positioning performance on real-world datasets, including an indoor environment with multi-floor structures.
+The tightly coupled IM-SLAM system achieves a final positioning error of less than 5 meters over an approximately 200-meter trajectory.
+  • We conduct ablation studies evaluating the impact of pose aiding for system initialization, barometric aiding for height stabilization, and inertial sensor quality.
+These analyses serve to clearly illustrate the overall performance bounds of the proposed systems.
 Reproducible research: Both the datasets and algorithms used in this study are available at https://github.com/Huang-Chuan/IM-SLAM.
 2 Preliminaries
-As magnetic field modeling is essential for magnetic field SLAM, we briefly review two commonly used magnetic field models in this section; the first model is for modeling the magnetic field at local small scales, and the second is at global large scales.
-Let us start with magnetostatics. Assuming that there is no time-varying electric field and current density in the region of interest, the magnetic field fulfills the curl-free condition [25, p. 180]
+As magnetic field modeling is essential for magnetic field SLAM, we briefly review two commonly used magnetic field models: the first captures local, small-scale variations, while the second describes global, large-scale fields.
+In the absence of time-varying electric fields and current density in the region of interest, the magnetic field fulfills the curl-free condition [25, p. 180]
 $\nabla\times M(r)=0$ (1)
@@
 $M(r)=\nabla\phi(r).$ (2)
-Both of the two magnetic field models presented next are based on (2).
+Both magnetic field models presented next are based on (2).
 2.1 Local Magnetic Field Model
@@
 $\phi(r;c)=\;\sum_{\begin{subarray}{c}i,j,k\geq 0\\ i+j+k\leq n+1\end{subarray}}c_{i,j,k}\,r_{x}^{i}r_{y}^{j}r_{z}^{k}.$ (3)
-Here $c_{i,j,k}\in\mathbb{R}$ denotes the coefficient and $r\triangleq[r_{x}\;r_{y}\;r_{z}]^{\top}$.
+Here $c_{i,j,k}\in\mathbb{R}$ denotes the coefficient and $r\triangleq[r_{x}\;r_{y}\;r_{z}]^{\top}$ denotes the position.
 Furthermore, the coefficients of the resulting magnetic field model $M(r;c)$ from applying (2) are further constrained by the divergence-free condition [25, p.
@@
 A global magnetic field model refers to a model that describes the magnetic field across the entire environment, such as a building or a large lab.
-The approximated GPmodel proposed in [23] allows for efficient regression with much less computational complexity than the full GPmodel.
+The approximated GP model proposed in [23] allows for efficient regression with much less computational complexity than a full GP model.
 The development of the magnetic field model starts with modeling the scalar magnetic potential function $\phi(r)$ as a GP
@@
 $\kappa_{\text{lin.}}(r,r^{\prime})$ $=\sigma_{\text{lin.}}^{2}r^{\top}r^{\prime}$ (7b)
-and
-$\kappa_{\text{SE}}(r,r^{\prime})=\sigma_{\text{SE}}^{2}\text{exp}\left(-\frac{\lVert r-r^{\prime}\rVert^{2}}{2l^{2}_{\text{SE}}}\right).$ (7c)
-Here $\sigma_{\text{lin.}}$ denotes the magnitude scale of the linear kernel $\kappa_{\text{lin.}}$, $\sigma_{\text{SE}}$ and $l_{\text{SE}}$ denote the magnitude scale and the characteristic length scale of the square exponential kernel $\kappa_{\text{SE}}$, respectively.
+$\kappa_{\text{SE}}(r,r^{\prime})$ $=\sigma_{\text{SE}}^{2}\exp\left(-\frac{\lVert r-r^{\prime}\rVert^{2}}{2l^{2}_{\text{SE}}}\right).$ (7c)
+Here $\mathcal{GP}\left(\mu,\kappa(x,x^{\prime})\right)$ denotes a collection of Gaussian random variables with mean $\mu$ and their covariance specified by $\kappa(x,x^{\prime})$.
+Furthermore, $\sigma_{\text{lin.}}$ denotes the magnitude scale of the linear kernel $\kappa_{\text{lin.}}$, $\sigma_{\text{SE}}$ and $l_{\text{SE}}$ denote the magnitude scale and the characteristic length scale of the square exponential kernel $\kappa_{\text{SE}}$, respectively.
 To reduce the computation complexity, the square exponential kernel is approximated by
@@
 where $m$ is the number of basic functions and $S_{\text{SE}}(\cdot)$ denotes the spectral density of $\kappa_{\text{SE}}$.
-Furthermore, $\psi_{j}(\cdot)$ and $\lambda_{j}$ denote the $j$-th eigenfunction and eigenvalue of the Laplace operator in the domain of a cuboid ($L_{x},L_{x}]\times[-L_{y},L_{y}]\times[-L_{z},L_{z}]$) with Dirichlet boundary conditions, respectively.
+Furthermore, $\psi_{j}(\cdot)$ and $\lambda_{j}$ denote the $j$-th eigenfunction and eigenvalue of the Laplace operator in the domain of a cuboid ($L_{x},L_{x}]\times[-L_{y},L_{y}]\times[-L_{z},L_{z}]$) with Dirichlet boundary conditions.
 They are given by
@@
 $\lambda_{j}^{2}$ $=\sum_{d=x,y,z}\left(\frac{\pi n_{j,d}}{2L_{d}}\right)^{2}$ (8c)
-respectively, where $n_{j,d}\in\mathbb{Z}^{+}$. Combining (7a) and (8a) and relating to the weight space view [26, Ch. 2], the scalar potential function is then approximated by
+where $n_{j,d}\in\mathbb{Z}^{+}$. Combining (7a) and (8a) and relating to the weight space view [26, Ch. 2], the scalar potential function is then approximated by
 $\phi(r;\eta)\approx\Psi(r)\eta,\quad\eta\in\mathbb{R}^{3+m}$ (9a)
 where
-$\Psi(r)=[r^{\top}\;\psi_{1}(r)\;\psi_{2}(r)\;\cdots\;\psi_{m}(r)]$ (9b)
-$\eta\sim\mathcal{N}(0,\Lambda)$ (9c)
+$\Psi(r)=[r^{\top}\;\psi_{1}(r)\;\psi_{2}(r)\;\cdots\;\psi_{m}(r)],\;\eta\sim\mathcal{N}(0,\Lambda)$ (9b)
 and
-$\Lambda=\text{diag}\left(\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},S_{\text{SE}}(\lambda_{1}),S_{\text{SE}}(\lambda_{2}),\cdots,S_{\text{SE}}(\lambda_{m})\right).$ (9d)
+$\Lambda=\text{diag}\left(\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},\sigma_{\text{lin}}^{2},S_{\text{SE}}(\lambda_{1}),S_{\text{SE}}(\lambda_{2}),\cdots,S_{\text{SE}}(\lambda_{m})\right).$ (9c)
 Here $\mathcal{N}(0,\Lambda)$ denotes a Gaussian distributed random variable with mean 0 and covariance matrix $\Lambda$.
@@
 $M(r;\eta)=\nabla\phi(r;\eta)\approx\nabla\Psi(r)\eta.$ (10)
-2.3 Discusssion
-The polynomial model presented in Section 2.1 is well-suited for local modeling, due to its simplicity and efficiency.
-The first-degree polynomial model was used in [11] to model the magnetic field within the small area covered by the magnetometer array at each time step (see Fig.
-2).
-In this model, both the spatial coordinates and the magnetic field are expressed in the body frame of the magnetometer array.
-On the other hand, the approximated GPmodel presented in Section 2.2 is more appropriate for global modeling, as the sinusoidal basis functions can capture the magnetic field variations across a larger area.
-It is more computationally demanding than the first-degree polynomial model.
+2.3 Discussion
+The polynomial model presented in Section 2.1 is well-suited for local modeling, due to its simplicity and computational efficiency.
+The first-degree polynomial model was used in [11] to model the magnetic field within the small area covered by the magnetometer array at each time step.
+In this model, both the spatial coordinates and the magnetic field are typically expressed in the body frame of the magnetometer array.
+On the other hand, the approximated GP model presented in Section 2.2 is appropriate for global modeling, as the sinusoidal basis functions can capture the magnetic field variations across a larger area.
 Typically, several hundreds to thousands of basis functions are used to represent the magnetic field across an area of several hundred square meters [7].
-Since this model aims to describe the field over a global region, both the spatial coordinates and the magnetic field are expressed in the navigation frame.
-Figure 2: Coordinate frames used by the local and global magnetic field models.
-The local magnetic field model at each time step aligns with the body frame of the magnetometer array, e.g., $\mathcal{B}_{1},\mathcal{B}_{2}$ and $\mathcal{B}_{3}$, while the global magnetic field model uses the navigation frame denoted by $O$ in the middle of the plot.
-3 Magnetic Field-Based Positioning Systems
-In this section, we briefly review the system models in two magnetic field-based positioning systems: the first is the magnetic field SLAM system proposed in [7], and the second is the MAINSproposed in [11].
-These systems, illustrated in Fig.
-3a and Fig.
-3b, are the foundations of the proposed IM-SLAMsystems.
-Figure 3: (a) An overview of the magnetic field SLAM system [7].
-The system performs dead reckoning using visual odometry data and uses a global magnetic field model to register a single magnetometer’s measurements, which are used to correct position drift.
-(b) An overview of the MAINS.
-The system performs inertial navigation using IMUmeasurements and uses a local magnetic field model to register a magnetometer array’s measurements, which are used to correct velocity estimates.
+Therefore, this model is computationally costly.
+Since this model aims to describe the field over a global region, both the spatial coordinates and the magnetic field are typically expressed in the navigation frame.
+3 Magnetic-Field-Based Positioning Systems
+In this section, we briefly review the system models used in two magnetic field-based positioning systems.
+The first model is used by the magnetic field SLAM system in [7], and the second by MAINS in [11].
 3.1 The Magnetic Field SLAM System
-The state-of-the-art magnetic field SLAM system [7] uses low-drift visual odometry and magnetic field measurements from a single magnetometer.
-The basic idea is to perform dead-reckoning using the visual odometry data and use magnetic field measurements to correct the accumulated position drift.
-The state vector of the magnetic field SLAM at time step $k$ is defined as
-$x_{k}^{\text{\tiny{SLAM}}}\triangleq[p_{k}^{\top}\;q_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (11)
-where $p_{k}\in\mathbb{R}^{3}$ denotes the position, $q_{k}\in S^{3}$ denotes the orientation in quaternion representation, and $\eta_{k}\in\mathbb{R}^{3+m}$ denotes the coefficients of the approximated GPmagnetic field model.
+The state-of-the-art magnetic field SLAM system [7] uses low-drift odometry and magnetic field measurements from a single magnetometer.
+The state vector of the magnetic field SLAM at time step $k$ is defined as $x_{k}^{\text{\tiny{SLAM}}}\triangleq[p_{k}^{\top}\;q_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$, where $p_{k}\in\mathbb{R}^{3}$ denotes the position, $q_{k}\in S^{3}$ denotes the orientation in quaternion representation, and $\eta_{k}\in\mathbb{R}^{3+m}$ denotes the coefficients of the approximated GP magnetic field model.
 The system model of the magnetic field SLAM is
@@
 In addition, $\otimes$ denotes quaternion multiplication, and $R_{k}\in SO(3)$ denotes the rotation matrix corresponding to the quaternion $q_{k}$.
-3.2 The Magnetic Field-Aided Inertial Navigation System
-The MAINSproposed in [11] uses a low-cost IMUand magnetometer array (see Fig.
-4) to provide low-drift inertial navigation.
-The basic idea is to use inertial navigation system (INS)to estimate poses, and magnetic field measurements to correct velocity estimates.
-The state vector of MAINS at time step $k$ is defined as
+3.2 The Magnetic-Field-Aided Inertial Navigation System
+MAINS proposed in [11] uses a low-cost IMU and a magnetometer array to provide low-drift inertial navigation. The state vector of MAINS at time step $k$ is defined as
 $x_{k}^{\text{\tiny{mains}}}\triangleq[{x_{k}^{\text{\tiny{ins}}}}^{\top}\;\theta_{k}^{\top}]^{\top}$ (12a)
@@
 Here $f^{\text{\tiny{ins}}}(\cdot)$ denotes the inertial subsystem dynamics, $f^{\theta}(\cdot)$ denotes the magnetic subsystem function.
-The exact form of these functions is presented in [11].
+The exact form of these functions are presented in [11].
 Moreover, $\tilde{u}_{k}=[\tilde{a}_{k}^{\top}\;\tilde{\omega}_{k}^{\top}]^{\top}$ denotes the IMU measurements, where $\tilde{a}_{k}\in\mathbb{R}^{3}$ and $\tilde{\omega}_{k}\in\mathbb{R}^{3}$ denote the specific force and angular velocity, respectively.
@@
 Finally, $e_{k}\in\mathbb{R}^{3N}$ denotes the white Gaussian measurement noise.
-Figure 4: The sensor board used in the magnetic field-aided inertial navigation system.
+Figure 1: The sensor board used in MAINS.
 It has 30 PNI RM3100 magnetometers and an Osmium MIMU 4844 IMU mounted on the bottom side.
@@
 Both the magnetic field SLAM system and the MAINS have their own advantages and limitations.
-The magnetic field SLAM system has a bounded positioning error in mapped regions.
-However, its reliance on the low-drift visual odometry limits its usability.
-The MAINS, on the other hand, can provide low-drift inertial navigation, but it only maintains a local magnetic field map and, thus, cannot provide positioning with bounded error.
-In the next section, we propose a loosely-coupled magnetic field SLAM system that combines the advantages of both systems while mitigating their limitations.
-4 The Loosely-Coupled IM-SLAMSystem
-A loosely coupled IM-SLAM system can be regarded as a cascaded architecture composed of a modified MAINSand a modified magnetic field SLAM system, as illustrated in Fig.
-5.
-The dependence on low-drift visual odometry data is eliminated by replacing it with odometry provided by the modified MAINSsystem.
-Furthermore, the measurement model in the magnetic field SLAM system is adapted to support magnetometer arrays, rather than a single magnetometer, as well as an additional barometer.
-The barometer supplies altitude information, enabling the SLAM system to achieve improved 3D trajectory estimation.
-Figure 5: An overview of the loosely-coupled IM-SLAMsystem. The system is a cascade of a modified MAINS, an odometry data conversion module, and a modified magnetic field SLAM system.
-4.1 Computing Odometry Data with the MAINS
-The original MAINS [11] estimates the pose at the current timestep; therefore, it cannot be used directly to compute odometry data.
-One way to make the MAINS support odometry data computation is to augment the state vector of MAINSto include the complete state vector (or the pose) at the past and compute the difference between the pose at the pase and the current.
+The magnetic field SLAM system has a bounded positioning error when in previously mapped areas.
+However, its reliance on the low-drift odometry limits its usability.
+MAINS, on the other hand, can provide low-drift inertial navigation, but it only maintains a local magnetic field map; thus, positioning errors grow without bound.
+In the next section, we propose a loosely coupled magnetic field SLAM system that combines the advantages of both systems while mitigating their limitations.
+4 The Loosely Coupled IM-SLAM System
+The loosely coupled IM-SLAM system integrates a modified MAINS [11] with a modified magnetic field SLAM system [7] in a cascaded configuration.
+It operates across two distinct time scales, following a common paradigm in visual SLAM systems, i.e., a high-rate odometry process and a low-rate optimization process [31].
+In our implementation, the modified MAINS runs at the IMU rate and is indexed by $k$, whereas the modified magnetic field SLAM system operates at a rate downsampled by a factor of $K$ and is indexed by $n$.
+The relationship between the two time indices is given by $k=1+nK$.
+This design reduces the overall computational load, as the magnetic field SLAM subsystem is computationally intensive, typically having a state dimension an order of magnitude larger than that of the modified MAINS.
+4.1 Computing Odometry Data with MAINS
+The original MAINS [11] estimates only the current pose; therefore, it cannot be used directly to compute odometry data.
+One approach to enable odometry data computation in MAINS is to augment its state vector with a past state (or pose), and compute the relative motion between the past and current poses.
 A detailed explanation of this modification is presented next.
 The augmented state vector is defined as
-$x_{k}^{a}$ $\triangleq[x_{k}^{\text{\tiny{mains}}\top}\;x_{k}^{\text{\tiny{past}}\top}]^{\top}$ (15a)
-$x_{1}^{\text{\tiny{past}}}$ $\triangleq x_{1}^{\text{\tiny{mains}}}$ (15b)
-$x_{k}^{\text{\tiny{past}}}$ $\triangleq f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}}),\quad k\geq 2$ (15c)
-$f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})$ $=\begin{cases}x_{k-1}^{\text{\tiny{past}}}&nK+3\leq k\leq(n+1)K+1\\ x_{k-1}^{\text{\tiny{mains}}}&k=nK+2\end{cases}$ (15d)
-$n$ $=0,1,2,\cdots$ (15e)
-Here $K\in\mathbb{N}$ denotes the number of time steps within the time interval for which the odometry data is computed. The state-space model of the modified MAINSis then
-$x_{k+1}^{a}=f^{a}(x_{k}^{a},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{mains}}}(x_{k}^{\text{\tiny{mains}}},\tilde{u}_{k},w_{k})\\ f^{\text{\tiny{past}}}(x_{k-1}^{\text{\tiny{past}}},x_{k-1}^{\text{\tiny{mains}}})\end{bmatrix}$ (16a)
+$\mathbf{x}^{a}_{k}\triangleq\begin{bmatrix}\mathbf{x}^{\text{\tiny{mains}}}_{k}\\ \mathbf{x}^{\text{\tiny{past}}}_{k}\end{bmatrix},\qquad\mathbf{x}^{\text{\tiny{past}}}_{1}\triangleq\mathbf{x}^{\text{\tiny{mains}}}_{1}.$ (14a)
+Here $x_{k}^{\text{\tiny{past}}}$ denotes the past state at the start of the interval over which the odometry data is computed. The state dynamics is defined as
+$x^{a}_{k+1}$ $=f^{a}\!\left(x^{a}_{k},\tilde{u}_{k},w_{k}\right)$ (14b)
+$f^{a}\!\left(x^{a}_{k},\tilde{u}_{k},w_{k}\right)$ $\triangleq\begin{bmatrix}f^{\text{\tiny{mains}}}\!\left(x^{\text{\tiny{mains}}}_{k},\tilde{u}_{k},w_{k}\right)\\ \begin{cases}x^{\text{\tiny{mains}}}_{k},&\text{if }k\in\mathcal{K}_{\mathrm{upd}}\\ x^{\text{\tiny{past}}}_{k},&\text{otherwise}\end{cases}\end{bmatrix}$.
+(14c)
+Here, $\mathcal{K}_{\mathrm{upd}}=\{1,\,1+K,\,1+2K,\dots\}$ denotes the set of time steps at which odometry data are computed. The state-space model of the modified MAINS is then
+$x_{k+1}^{a}$ $=f^{a}(x_{k}^{a},\tilde{u}_{k},w_{k})$ (15a)
 $y_{k}$ $=H^{a}x_{k}^{a}+e_{k}$ (15b)
-$H^{a}$ $=\begin{bmatrix}0_{3\times 16}&\Phi(r^{(1)})&0_{3\times 16}\\ \vdots&\vdots&\vdots\\ 0_{3\times 16}&\Phi(r^{(N)})&0_{3\times 16}\end{bmatrix}$. (16c)
-The modified MAINS can compute odometric data between the time step $i=nK+1$ and $j=(n+1)K+1$ using the posterior state estimate at time step $j$ as
-$\tilde{u}^{o}_{i,j}$ $=\begin{bmatrix}\Delta p_{i,j}\\ \Delta q_{i,j}\end{bmatrix}=\begin{bmatrix}\hat{p}_{j}-\hat{p}_{i}\\ \hat{q}_{i}^{\ast}\otimes\hat{q}_{j}\end{bmatrix}$ (17a)
-where $(\hat{\cdot}_{i})$ and $(\hat{\cdot}_{j})$ denote the smoothed and filtered quantity at time $i$ and $j$, respectively, and $(\cdot)^{\ast}$ denotes conjugated quaternion.
+$H^{a}$ $=\begin{bmatrix}H^{\text{\tiny{mains}}}\quad 0\end{bmatrix}$. (15c)
+The error-state Kalman filter [27] is applied here for state estimation.
+To avoid the double use of the magnetometer measurements in the modified magnetic field SLAM system presented in the next subsection, no measurement update is performed at any time in $\mathcal{K}_{\mathrm{upd}}$.
+The modified MAINS computes the odometric data between the time step $i=1+nK$ and $j=1+(n+1)K$ using the state estimate at time step $j$ as
+$\tilde{u}^{o}_{n}$ $=\begin{bmatrix}\Delta p_{i,j}\\ \Delta q_{i,j}\end{bmatrix}=\begin{bmatrix}\hat{p}_{j}-\hat{p}_{i}\\ \hat{q}_{i}^{\ast}\otimes\hat{q}_{j}\end{bmatrix}$ (16a)
+where $(\hat{\cdot}_{i})$ and $(\hat{\cdot}_{j})$ denote the smoothed and filtered quantities at time $i$ and $j$, respectively, and $(\cdot)^{\ast}$ denotes the quaternion conjugation.
 The covariance of the odometry data can be approximated by
-$\text{Cov}(\tilde{u}_{i,j}^{o})\approx A_{i,j}\;\text{Cov}(\hat{p}_{i},\hat{p}_{j},\hat{q}_{i},\hat{q}_{j})\;A_{i,j}^{\top}$ (17b)
+$\text{Cov}(\tilde{u}_{n}^{o})\approx A_{i,j}\;\text{Cov}(\hat{p}_{i},\hat{p}_{j},\hat{q}_{i},\hat{q}_{j})\;A_{i,j}^{\top}$ (16b)
 $A_{i,j}=\begin{bmatrix}I_{3}&-I_{3}&0_{3}&0_{3}\\ 0_{3}&0_{3}&I_{3}&-\hat{R}_{i}^{\top}\hat{R}_{j}\end{bmatrix}$ (16c)
-where $\text{Cov}(\cdot)$ denotes the covariance. Note the covariance of the quaternion is defined in the tangent space of $S^{3}$, see [27, p. 54] for details.
+where $\text{Cov}(\cdot)$ denotes the covariance.
+Note that the covariance matrix on the right hand side of (16b) can be extracted from the covariance matrix computed by the filter, and the covariance of the quaternion is defined in the tangent space of $S^{3}$, see [27, p.
+54] for details.
 4.2 The Modified Magnetic Field SLAM System
-The modified magnetic field SLAM system uses the odometric data provided by the modified MAINS, magnetic field measurements from the sensor array, and barometer measurements.
+The modified magnetic field SLAM system uses the odometric data $\tilde{u}^{o}_{n}$ provided by the modified MAINS, magnetic field measurements from the sensor array at time steps in $\mathcal{K}_{\mathrm{upd}}$, and barometer measurements.
 The state vector of this system and the dynamics remain the same as in Section 3.1.
 The measurement equation is modified to
-$y_{k}=h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})+e_{k}$ (18a)
+$y_{n}=h^{\text{\tiny{array-slam}}}(x_{n}^{\text{\tiny{SLAM}}})+e_{n}$ (17a)
 where
-$h^{\text{\tiny{array-slam}}}(x_{k}^{\text{\tiny{SLAM}}})=\begin{bmatrix}R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(1)}\big)\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(2)}\big)\\ \vdots\\ R_{k}^{\top}\nabla\Psi\big(p_{k}+R_{k}r^{(N)}\big)\end{bmatrix}\eta_{k}.$ (18b)
+$h^{\text{\tiny{array-slam}}}(x_{n}^{\text{\tiny{SLAM}}})=\begin{bmatrix}R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(1)}\big)\\ R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(2)}\big)\\ \vdots\\ R_{n}^{\top}\nabla\Psi\big(p_{n}+R_{n}r^{(N)}\big)\end{bmatrix}\eta_{n}.$ (17b)
 Here $h^{\text{\tiny{array-slam}}}(\cdot)$ denotes the modified measurement function, which takes into account the multiple sensors and the geometric configuration of the sensor array.
-This measurement equation is augmented when barometric measurements are available, i.e.,
-$y_{k}^{b}=p_{z,k}+e_{k}^{b}$ (19)
-where $p_{z,k}\in\mathbb{R}$ denotes the altitude (the third component of the position vector), and $e_{k}^{b}\in\mathbb{R}$ denotes the barometer measurement noise, modeled as Gaussian white noise.
+This measurement equation is augmented when barometric measurements are available at time steps in $\mathcal{K}_{\mathrm{upd}}$, i.e., $y_{n}^{b}=p_{z,n}+e_{n}^{b}$, where $p_{z,n}\in\mathbb{R}$ denotes the altitude (the third component of the position vector), and $e_{n}^{b}\in\mathbb{R}$ denotes the barometer measurement noise, modeled as Gaussian white noise.
 4.3 Summary
-The loosely-coupled IM-SLAM system combines the advantages of both the magnetic field SLAM system and the MAINS.
-It is a self-contained system that can provide bounded-error positioning in mapped areas without the need for external low-drift odometry.
-However, there are two main limitations of this system.
-First, the magnetometer measurements are used twice: first in the modified MAINSand then in the modified magnetic field SLAM system.
-Second, the odometric data provided by the modified MAINSis correlated over time, which violates the assumption of uncorrelated odometric data in the magnetic field SLAM system.
-These two limitations can potentially degrade the performance of the loosely-coupled IM-SLAM system.
-To address these limitations, we propose a tightly-coupled IM-SLAM system in the next section.
-5 The Tightly-Coupled IM-SLAMSystem
-The tightly-coupled IM-SLAM system integrates the inertial, barometric, and magnetic sensors into a unified SLAM framework (see Fig.
-6).
-In this framework, the state vector encompasses the full set of states, including the inertial states, local and global magnetic field model coefficients.
+The loosely coupled IM-SLAM system is a straightforward implementation based on the magnetic field SLAM system and MAINS.
+However, the main issue is that the odometric data provided by the modified MAINS is correlated over time, which violates the assumption of uncorrelated odometric data in the magnetic field SLAM system.
+Notably, the temporal correlations in the odometry data are not unique to the proposed system.
+Visual and visual–inertial odometry estimates are subject to the same effect, since successive motion estimates are computed from overlapping measurements [29].
+Nevertheless, this limitation can potentially degrade the performance of the loosely coupled IM-SLAM system.
+To address this limitation, we propose a tightly coupled IM-SLAM system in the next section.
+5 The Tightly Coupled IM-SLAM System
+The tightly coupled IM-SLAM system integrates the inertial, barometric, and magnetic sensors into a unified SLAM framework.
+In this framework, the state vector encompasses the full set of states, including the inertial navigation states, and the local and global magnetic field model coefficients.
 The system dynamics and measurement equation are jointly formulated to leverage all available information to achieve better performance.
-In the following, the state-space model of this system is presented.
-Figure 6: An overview of the tightly-coupled IM-SLAMsystem.
-This system consists of an inertial navigation system, a barometer model, and a local and global magnetic field model.
-These components are integrated into a unified SLAMframework.
+Next, the state-space model of this system is presented.
 5.1 State Dynamics
-The state vector of the tightly-coupled SLAM system at time step $k$ is defined as
-$x_{k}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}\;\theta_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$ (20)
-The state transition model is
-$x_{k+1}=f(x_{k},\tilde{u}_{k},w_{k})$ (21)
+The state vector of the tightly coupled SLAM system at time step $k$ is defined as $x_{k}\triangleq[p_{k}^{\top}\;v_{k}^{\top}\;q_{k}^{\top}\;b_{a,k}^{\top}\;b_{g,k}^{\top}\;\theta_{k}^{\top}\;\eta_{k}^{\top}]^{\top}$, and the state transition model is
+$x_{k+1}=f(x_{k},\tilde{u}_{k},w_{k})$ (18a)
 where
 $f(x_{k},\tilde{u}_{k},w_{k})=\begin{bmatrix}f^{\text{\tiny{ins}}}(x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\text{\tiny{ins}}}_{k})\\ f^{\theta}(\theta_{k},x^{\text{\tiny{ins}}}_{k},\tilde{u}_{k},w^{\theta}_{k})\\ \eta_{k}\end{bmatrix}$.
-(22)
+(18b)
 5.2 Measurement Equations
 Since the state vector includes both the local and global magnetic field model coefficients, the magnetometer array’s measurements can be expressed in terms of both models.
-To achieve IM-SLAM, one can use the measurement equations in an alternating manner from either the MAINSor the magnetic field SLAM system.
+To exploit their complementary advantages, one can use the measurement equations in the MAINS or the magnetic field SLAM system in an alternating manner.
 A natural choice is to use the computationally efficient measurement equation from the MAINS most of the time, while occasionally using the more complex measurement equation from the magnetic field SLAM system.
 In this way, the system can benefit from the low-drift inertial navigation provided by the MAINS while still leveraging the global consistency of the magnetic field SLAM system.
-In the proposed tightly-coupled IM-SLAM, the measurement equation of the global model, which is based on (18), is further simplified to reduce computational complexity.
-The basic idea is that the magnetic field at any sensor’s location can be viewed as a superimposition of the magnetic field in the center of the array and some local magnetic field variation.
+In the proposed tightly coupled IM-SLAM, the measurement equation of the global model, which is based on (17), is further simplified to reduce computational complexity.
+The basic idea is that the magnetic field at any sensor’s location can be viewed as a superposition of the magnetic field in the center of the array and some local magnetic field variation.
 This results in
@@
 Furthermore, $\text{grad}(\theta_{k})\in\mathbb{R}^{3\times 3}$ has a clear physical interpretation, i.e., the gradient of the local magnetic field.
-To sum up, there are two measurement equations for magnetometer array measurements: the equation used by the MAINSand the equation (23).
-They are used in an alternating manner: the former is employed at every update step except once $D$ samples, when the second model is used instead.
+To sum up, there are two measurement equations for magnetometer array measurements: the equation (15c) used by MAINS and the equation (19).
+They are used in an alternating manner: the former is employed at every update step except once every $D$ samples, when the second model is used instead.
 That is,
@@
 Lastly, similar to what was done for the loosely coupled IM-SLAM system, the measurement equation is augmented when barometer measurements are available.
+5.3 Summary
+The tightly coupled IM-SLAM system augments the INS state with both local and global magnetic field model coefficients, performing measurement updates with these models in an alternating fashion.
+By estimating the navigation state in a single step, the system bypasses the odometry data computation described in 4.1, thereby resolving the temporal correlation issues inherent in loosely coupled IM-SLAM.
+However, the computational demand is significantly higher; the state dimension is comparable to that of the modified magnetic field SLAM system and operates at the IMU rate, even though the complex measurement update in (19) is invoked less frequently.
 6 Experimental Results
-To evaluate the performance of the proposed loosely-coupled and tightly-coupled IM-SLAMsystems, we conducted experiments in an indoor environment using a custom-built sensor array, see Fig.
-4.
-The sensor array is equipped with an IMU, a barometer, and an array of magnetometers.
+To evaluate the performance of the proposed loosely coupled and tightly coupled IM-SLAM systems, we conducted experiments in an indoor environment using a custom-built sensor array, see Fig.
+1.
+The sensor array is equipped with an IMU, a barometer, and 30 magnetometers.
 In each experiment, a person carried the sensor array and started moving in a room equipped with a motion capture system for several tens of seconds.
@@
 In total, three types of trajectories were recorded: a trajectory with long straight segments, a trajectory with square loops, and a trajectory with spiral upward segments (climbing a spiral staircase).
-The characteristics of the datasets are summarized in TABLE 1.
+They are named Long Corridor (LC), Corridor (C), and Spiral Staircase (SS).
+The characteristics of these datasets are summarized in TABLE 1.
 The IMU and magnetometers were sampled at 100 Hz, while the barometer was sampled at 10 Hz.
 When available, the motion capture system provided ground truth poses at 100 Hz.
-In the loosely-coupled IM-SLAM system, the odometry data was internally computed at $5$ Hz (K=20K\!=\!20).
-In the tightly-coupled SLAM system, the magnetic field model switching time was set to 1 second (D=100D\!=\!100), the length scale of the approximated GPmodel $l_{\text{SE}}$ was set to 0.7 meters.
-More details about the parameter settings can be found in the code repository.
-The standard error-state extended Kalman filter (EKF)is used in all systems for state estimation; interested reader are referred to [27] for details.
+In the loosely coupled IM-SLAM system, the odometry data was internally computed every second (K=100K\!=\!100).
+In the tightly coupled SLAM system, the magnetic field model switching time was set to 1 second (D=100D\!=\!100), the length scale of the approximated GP model $l_{\text{SE}}$ was set to 0.7 to 1 meter depending on the datasets.
+The choice of $D$ and $K$ was based on experimental results.
+Too large a value would cause SLAM systems to degrade to MAINS, while too small a value would cause the filter to diverge.
+A study on the effect of model switching time is conducted in Section 6.4.
 Table 1: Information about the datasets
@@
 […]
+[…]
+[…]
+[…]
+[…]
 * including the initial part of the trajectory in the motion capture area, $\dagger$ excluding the initial part of the trajectory in the motion capture area.
@@
 LC: long corridor   C: corridor   SS: spiral staircase.
-6.1 Comparison with the MAINS
-In the first experiment, the loosely-coupled and tightly-coupled IM-SLAM systems were compared with MAINS.
+6.1 Comparison with MAINS
+In the first experiment, the loosely coupled and tightly coupled IM-SLAM systems were compared with MAINS.
 Note that MAINS was modified in the same way as in Section 3.1 to integrate barometer measurements.
-All three systems used the pose from the motion capture system as additional measurements when the sensor array was in the room at the beginning of the trajectory.
-This is to stablize the IMU bias estimates, which is important for good positioning performance.
+All three systems used the pose from the motion capture system as additional measurements at the beginning of the trajectory.
+This is to stabilize the IMU bias estimates, which is important for good positioning performance [30, Ch.
+22].
+For many applications, such as first-responder localization, one can expect to have similar aiding at the start, such as GPS positioning.
 During the later parts of the trajectory, the systems relied solely on measurements from the onboard sensors for positioning.
 Three examples of the trajectories estimated by the three systems are shown in Fig.
-7, Fig.
-8, and Fig.
-9, respectively.
-The quantitative results are summarized in TABLE 2 and TABLE 3, where the horizontal error, vertical error, and yaw error at the end of the trajectories are reported.
+2, Fig.
+3, and Fig.
+4, respectively.
+The quantitative results are summarized in TABLE 2, where the horizontal error and vertical error at the end of the trajectories are reported.
 The horizontal error is defined as the Euclidean distance between the estimated position and the ground truth position projected onto the horizontal plane, while the vertical error is defined as the absolute difference between the estimated altitude and the ground truth altitude.
-The yaw error is defined as the absolute difference between the estimated yaw angle and the ground truth yaw angle.
 It can be seen from the figures that all systems can track the general shape of the trajectories.
-The noticeable difference between the tightly coupled IM-SLAMand other systems is that the estimated trajectories by the former have more jerky segments, which occur when the systems switch between the two measurement models; see (25).
-This phenomenon could be attributed to two different modeling of the magnetic field models: the model in the first row of (25) is parameterized with the position of the body frame in the navigation frame while the model in the second row is parameterized with the (local) positions of the magnetometers in the body frame.
-This implies that the former allows direct observation of the position while the latter does not [28].
-The sudden jump observed on the estimated trajectory is a sign of gaining positional information when switching from the latter to the former model.
-Another observation is that the trajectories estimated by the SLAM systems during the long straight segments are more concentrated than those estimated by MAINS (see Fig.
-7), which indicates that the SLAM systems can better mitigate the drift during these segments.
-Moreover, from TABLE 2 and TABLE 3, it can be seen that the tightly-coupled IM-SLAM system outperforms both MAINS and the loosely-coupled IM-SLAM system in most cases.
-Table 2: Horizontal (Vertical) Error at the end of the Trajectories with a Barometer. Unit: meter
+The noticeable difference between the IM-SLAM systems and MAINS is that the estimated trajectories by the former have more jerky segments, which occur when the loosely coupled system uses (17) and the tightly coupled system uses (19) to update position estimates.
+These jerky segments are signs of gaining absolute positional information because the measurement equations (17) and (19) are parameterized with the absolute position of the body frame, while (15c) does not allow observation of absolute positions [28].
+One way to smooth the trajectory is to perform an extended Kalman filter (EKF) smoothing procedure following the measurement updates concerning (17) and (19).
+For the tightly coupled system, one can also introduce a pseudo measurement and perform measurement update together with (19), see Section 6.5.
+Another observation is that the loosely coupled system and the tightly coupled system performed noticeably differently on Dataset Long corridor-2, where a trajectory with two long segments (branches) and a trajectory with two overlapping segments are reported by them respectively.
+Given that similar magnetic field magnitudes were observed on these two branches of the trajectory (the loosely coupled system), it is very likely that they were, in fact, overlapping, indicating that the tightly coupled system performs better in this case.
+Moreover, from TABLE 2, it can be seen that the tightly coupled IM-SLAM system outperforms both MAINS and the loosely coupled IM-SLAM system in most cases.
+Table 2: Horizontal and Vertical Error at the End of the Trajectories with a Barometer. Unit: meter
 […]
-Long corridor-1 3.53 (0.29) 9.20 (0.91) 0.90 (0.22)
-Long corridor-2 2.28 (0.17) 5.23 (1.01) 1.51 (0.16)
-Long corridor-3 1.25 (0.13) 2.44 (0.42) 1.08 (0.10)
-Corridor-1 5.22 (0.14) 10.46 (0.27) 2.57 (0.14)
-Corridor-2 1.93 (0.09) 1.03 (0.77) 1.93 (0.13)
-Corridor-3 1.98 (0.03) 1.76 (0.10) 1.78 (0.05)
-Spiral staircase-1 3.55 (0.02) 5.22 (0.26) 1.79 (0.00)
-Spiral staircase-2 3.59 (0.17) 3.45 (0.12) 2.43 (0.16)
-Spiral staircase-3 8.37 (0.20) 9.13 (0.84) 3.20 (0.21)
+Long corridor-1 3.53 | 0.29 0.22 | 0.84 0.90 | 0.22
+Long corridor-2 2.28 | 0.17 1.52 | 0.60 1.51 | 0.16
+Long corridor-3 1.25 | 0.13 0.85 | 0.42 1.08 | 0.10
+Corridor-1 5.22 | 0.14 9.39 | 0.27 2.57 | 0.14
+Corridor-2 1.93 | 0.09 1.25 | 0.77 1.93 | 0.13
+Corridor-3 1.98 | 0.03 1.77 | 0.09 1.78 | 0.05
+Spiral staircase-1 3.55 | 0.02 3.46 | 0.19 1.79 | 0.00
+Spiral staircase-2 3.59 | 0.17 3.21 | 0.15 2.43 | 0.16
+Spiral staircase-3 8.37 | 0.20 10.00 | 0.72 3.20 | 0.21
 *: loosely coupled IM-SLAM system $\dagger$: tightly coupled IM-SLAM system
-Table 3: Yaw Error at the end of the Trajectories with a Barometer. Unit: degree
-[…]
-Long corridor-1 1.45 1.59 0.28
-Long corridor-2 0.79 1.68 2.34
-Long corridor-3 0.29 1.30 1.65
-Corridor-1 2.02 2.85 1.77
-Corridor-2 3.00 1.38 1.05
-Corridor-3 2.19 1.57 0.52
-Spiral staircase-1 2.32 1.06 1.83
-Spiral staircase-2 0.64 2.35 1.40
-Spiral staircase-3 1.58 2.49 0.52
-*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
-Figure 7: Trajectories (Long corridor-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
+Figure 2: Trajectories (Long corridor-2) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system.
 The room where the motion capture system is located is marked with a blue rectangle frame.
-Figure 8: Trajectories (Corridor-3) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
+Figure 3: Trajectories (Corridor-3) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system.
 The room where the motion capture system is located is marked with a blue rectangle frame.
-Figure 9: Trajectories (Spiral staircase-2) estimated by (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
+Figure 4: Trajectories (Spiral staircase-2) estimated by (a) MAINS, (b) the loosely coupled IM-SLAM system, and (c) the tightly coupled IM-SLAM system.
 The room where the motion capture system is located is marked with a blue rectangle frame.
 6.2 Ablation Study on the Effect of Barometer and Inertial Sensor Quality
-In the second part of the experiment, the effect of the barometric measurements and inertial sensor quality on the performance of the algorithms was investigated.
-For the first part, the barometer measurements were removed from all three systems, and the results are summarized in TABLE 4 and TABLE 5.
-Comparing TABLE 4 with TABLE 2, it can be seen that both the vertical error and horizontal error of all three systems increase slightly without the barometer measurements, but the overall performance remains unchanged.
-Furthermore, the yaw error of all three systems do not change significantly as seen from TABLE 5.
+In the second part of the experiment, the effect of the barometric measurements and inertial sensor quality on the performance of all systems was investigated.
+For the first part, the barometer measurements were removed from all three systems, and the results are summarized in TABLE 3.
+Comparing TABLE 3 with TABLE 2, it can be seen that both the vertical and horizontal errors of all three systems changed slightly (except the loosely coupled IM-SLAM on Long corridor-1) without the barometer measurements, but the overall performance remains unchanged.
 This indicates that the IM-SLAM systems are not heavily reliant on the barometer measurements.
 However, the barometer measurements are expected to have a more significant impact if explicit loop closure detection is to be performed in the SLAM systems, since the altitude information can help to distinguish different floors in a multi-floor building.
-For the second part, the IMU measurements quality is varied by averaging the measurements from different numbers of IMU sensors in Osmium MIMU 4844 IMU.
+For the second part, the IMU measurement quality is varied by averaging the measurements from different numbers of IMU sensors in Osmium MIMU 4844 IMU.
 The box plots of the horizontal errors at the end of the trajectories vs.
 the number of IMU sensors are shown in Fig.
-10, 11, and 12.
-It can be seen that the horizontal errors of the tightly-coupled IM-SLAM system are less sensitive to the number of IMU sensors compared to the other two systems, and it outperforms the other two algorithms on the datasets Long corridor and Corridor.
-On Spiral staircase datasets, the performance of the tightly-coupled IM-SLAM system improves as the number of IMU sensors increases, but it is not as good as MAINSwhen the number of IMU sensors is less than 4.
-We do not have a clear explanation for this phenomenon yet, but it could be related to the fact that the datasets Spiral staircase have more dynamic motions (e.g., turning and climbing) than the other two types of datasets.
-Further investigation is needed to fully understand this phenomenon.
-Lastly, the sudden increase in horizontal error when using 32 IMUsensors is due to exceptionally worse performance on one of the datasets (Spiral staircase-3), which is currently under investigation.
-Table 4: Horizontal (Vertical) Error at the End of the Trajectories without a Barometer. Unit: meter
+5.
+It can be seen that the horizontal errors of the tightly coupled IM-SLAM system are less sensitive to the number of IMU sensors compared to the other two systems.
+Table 3: Horizontal and Vertical Error at the End of the Trajectories without a Barometer. Unit: meter
 […]
-Long corridor-1 3.80 (1.01) 6.84 (6.75) 0.97 (0.83)
-Long corridor-2 2.33 (1.16) 8.41 (1.21) 1.49 (0.11)
-Long corridor-3 1.19 (1.21) 4.30 (0.89) 1.23 (0.43)
-Corridor-1 5.35 (0.54) 11.05 (0.36) 2.45 (0.28)
-Corridor-2 1.82 (0.44) 0.99 (0.95) 1.76 (0.32)
-Corridor-3 2.02 (0.52) 1.80 (0.36) 1.89 (0.51)
-Spiral staircase-1 3.00 (0.41) 6.90 (1.08) 1.74 (0.16)
-Spiral staircase-2 3.39 (0.10) 4.33 (0.90) 2.51 (0.22)
-Spiral staircase-3 8.22 (0.65) 8.48 (2.14) 4.13 (0.91)
+Long corridor-1 3.80 | 1.01 12.80 | 5.23 0.97 | 0.83
+Long corridor-2 2.33 | 1.16 1.07 | 0.82 1.49 | 0.11
+Long corridor-3 1.19 | 1.21 1.02 | 0.54 1.23 | 0.43
+Corridor-1 5.35 | 0.54 10.53 | 0.42 2.45 | 0.28
+Corridor-2 1.82 | 0.44 1.20 | 0.95 1.76 | 0.32
+Corridor-3 2.02 | 0.52 1.80 | 0.38 1.89 | 0.51
+Spiral staircase-1 3.00 | 0.41 3.85 | 0.93 1.74 | 0.16
+Spiral staircase-2 3.39 | 0.10 1.77 | 0.49 2.51 | 0.22
+Spiral staircase-3 8.22 | 0.65 9.00 | 2.44 4.13 | 0.91
 *: loosely coupled IM-SLAM system $\dagger$: tightly coupled IM-SLAM system
-Table 5: Yaw Error at the End of the Trajectories without a Barometer. Unit: degree
-[…]
-Long corridor-1 0.87 2.85 0.29
-Long corridor-2 1.24 2.17 2.31
-Long corridor-3 0.23 1.76 2.00
-Corridor-1 2.11 3.12 1.32
-Corridor-2 2.90 1.37 0.90
-Corridor-3 2.11 1.60 0.28
-Spiral staircase-1 2.25 0.23 1.62
-Spiral staircase-2 0.56 2.20 0.80
-Spiral staircase-3 1.71 0.41 1.40
-*: loosely coupled IM-SLAMsystem $\dagger$: tightly coupled IM-SLAMsystem
-Figure 10: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Long corridor datasets.
-The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
-Figure 11: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Corridor datasets.
-The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
-Figure 12: Box plot of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors on the Spiral staircase datasets.
-The algorithms from left to right are (a) MAINS, (b) the loosely-coupled IM-SLAM system, and (c) the tightly-coupled IM-SLAM system.
+(a) Dataset: Long corridor
+(b) Dataset: Corridor
+(c) Dataset: Spiral staircase
+Figure 5: Box plots of horizontal errors (red: median) at the end of trajectories of all algorithms using different numbers of IMU sensors.
+The box plots in each subfigure from left to right correspond to MAINS, the loosely coupled IM-SLAM system, and the tightly coupled IM-SLAM system.
+6.3 Ablation Study on the Effect of Pose Aiding Duration on Positioning Performance
+In the third part of the experiment, the effect of pose aiding duration on positioning performance was investigated.
+The pose aiding duration at the start was set from 0 to 30 seconds, and the horizontal error at the end of the estimated trajectory is calculated, as shown in TABLE 4.
+It can be seen that all systems have reasonably good positioning performance when the pose aiding duration is more than 20 seconds.
+Therefore, it is recommended to initialize these systems with at least 20 seconds of pose aiding before use.
+Table 4: Horizontal Error at the End of the Trajectories with Different Pose Aiding Time. Unit: meter
+Dataset Algorithm 0 s 10 s 20 s 30 s
+Corridor-3 MAINS 4.66 1.33 2.05 0.95
+IM-SLAM (L.) 6.36 1.99 3.65 1.22
+IM-SLAM (T.) 3.12 4.12 3.52 1.62
+Long corridor-2 MAINS 10.72 7.51 3.56 2.28
+IM-SLAM (L.) 9.03 6.66 1.79 0.87
+IM-SLAM (T.) 4.39 0.58 0.57 1.51
+Spiral staircase-2 MAINS 1.29 0.93 1.45 2.39
+IM-SLAM (L.) 1.70 0.96 1.26 1.30
+IM-SLAM (T.) 3.74 3.46 5.11 2.39
+6.4 Study on the Effect of Model Switching Time
+In the fourth part of the experiment, the effect of model switching time is studied.
+The value of $D$ is varied from 5 to 1500 and $\infty$, corresponding to model switching times of 0.05 s, 15 s, and no switching, respectively.
+The resulting horizontal error at the end of the estimated trajectory is reported in Table 5.
+It can be seen that when the switching time is 1 second, the tightly coupled SLAM system has the best overall positioning performance.
+Given the walking speed (approximately 1 m/s), the distance between two consecutive measurement updates concerning (19) is approximately 1 m, which is roughly on the same order of the length scale of the approximated GP model (0.7 to 1 m).
+As magnetic fields at locations one length scale apart are less correlated, the tightly coupled SLAM system is very likely to have “recorded” magnetic field measurements (without affecting navigation states estimation) in the exploration phase and used this information to correct drift when revisiting the same locations.
+Table 5: Horizontal Error at the End of the Trajectories with Different Model Switching Time. Unit: meter
+Dataset 0.05 s 1 s 5 s 15 s $\infty$ s
+Long corridor-1 27.70 0.9 5.52 1.28 3.53
+Long corridor-2 7.43 1.51 1.16 1.10 2.28
+Long corridor-3 20.70 1.08 1.30 1.10 1.25
+Corridor-1 1.87 2.57 2.60 5.83 5.22
+Corridor-2 10.11 1.93 2.71 1.89 1.93
+Corridor-3 18.61 1.78 0.69 3.06 1.98
+Spiral staircase-1 5.48 1.79 1.41 3.55 3.55
+Spiral staircase-2 9.42 2.43 2.91 2.53 3.59
+Spiral staircase-3 8.02 3.20 3.73 7.99 8.37
+6.5 Introducing Pseudo Measurements
+In the last part of the experiments, we investigate whether a pseudo measurement can be used to smooth the trajectories in the tightly coupled SLAM system.
+The measurement equation (19) is augmented with a pseudo measurement,
+$y^{\text{\tiny{pseudo}}}_{k}$ $=h^{\text{\tiny{pseudo}}}(x_{k})+e_{k}^{\text{\tiny{pseudo}}}$ (22a)
+$h^{\text{\tiny{pseudo}}}(x_{k})$ $=R^{\top}_{k}\nabla\Psi(p_{k})\eta_{k}-\Phi(0)\theta_{k}.$ (22b)
+The pseudo measurement is the difference between magnetic fields at the center of the magnetometer array predicted the Gaussian process model and the polynomial model.
+Ideally, the pseudo measurement should be zero, and the covariance of $e_{k}^{\text{\tiny{pseudo}}}$ reflects the strictness of this constraint.
+This pseudo measurement was introduced in the experiments on the datasets Corridor-3 and Spiral staircase-2.
+The dataset Long corridor was not selected for this experiment because the trajectory in Figure 2 was less jerky.
+The covariance of $e_{k}^{\text{\tiny{pseudo}}}$ was set to $0.005I_{3}$ and $0.01I_{3}$, respectively.
+Here, $I_{3}$ denotes the identity matrix.
+The trajectories reported by the tightly coupled IM-SLAM system after introducing the pseudo measurement are shown in Figure 6.
+Comparing the trajectories in Figure 3-(c) and Figure 4-(c), it can be seen that the trajectories in Figure 6 are smoother.
+Figure 6: These two figures (a) Trajectory (Corridor-3) (b) Trajectory (Spiral Staircase-2) illustrate the smoothing effect after introducing the pseudo measurement in the tightly coupled IM-SLAM system.
 7 Conclusion
-In this paper, a loosely coupled and a tightly coupled IM-SLAMsystem were proposed.
-Both systems rely solely on commonly available low-cost sensors: an inertial measurement unit (IMU), a magnetometer array, and a barometer.
-The proposed systems have position errors on the order of meters per 100 meters traveled, which showcases the feasibility of developing magnetic field-based SLAMsystems with low-cost sensors.
+In this paper, loosely coupled and tightly coupled IM-SLAM systems are proposed.
+Both systems rely solely on commonly available low-cost sensors: an inertial measurement unit, 30 magnetometers, and a barometer.
+The proposed systems have position errors on the order of meters per 100 meters traveled, which showcases the feasibility of developing magnetic-field-based SLAM systems with low-cost sensors.
 Experimental results demonstrate that the tightly coupled IM-SLAM system achieves better performance than the loosely coupled counterpart.
-This improvement arises because the tightly coupled system integrates the local and global magnetic field model within a unified framework, rather than employing them in a two-stage manner as in the loosely coupled design.
-Furthermore, the experiments show that incorporating a barometer primarily benefits IM-SLAMperformance in the vertical direction, and that the IMU quality can be degraded by approximately a factor of 32 without substantially affecting the overall accuracy.
-These results demonstrate the potential of applying the proposed systems in emergency response scenarios such as mine/fire rescue.
+Furthermore, the experiments show that incorporating a barometer primarily benefits IM-SLAM performance in the vertical direction, and that the IMU quality can be degraded by approximately a factor of 32 without substantially affecting the overall accuracy of the tightly coupled SLAM system.
+These results demonstrate the potential of applying the proposed systems in emergency response scenarios such as mine or fire rescue, where one cannot rely on a GNSS or visual-based localization system.